### v1

# Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics

Abstract
Low-level 3D representations, such as point clouds, meshes, NeRFs, and 3D Gaussians, are commonly used to represent 3D objects or scenes. However, humans usually perceive 3D objects or scenes at a higher level as a composition of parts or structures rather than points or voxels. Representing 3D as semantic parts can benefit further understanding and applications. We aim to solve part-aware 3D reconstruction, which parses objects or scenes into semantic parts. In this paper, we introduce a hybrid representation of superquadrics and 2D Gaussians, trying to dig 3D structural clues from multi-view image inputs. Accurate structured geometry reconstruction and high-quality rendering are achieved at the same time. We incorporate parametric superquadrics in mesh forms into 2D Gaussians by attaching Gaussian centers to faces in meshes. During the training, superquadrics parameters are iteratively optimized, and Gaussians are deformed accordingly, resulting in an efficient hybrid representation. On the one hand, this hybrid representation inherits the advantage of superquadrics to represent different shape primitives, supporting flexible part decomposition of scenes. On the other hand, 2D Gaussians are incorporated to model the complex texture and geometry details, ensuring high-quality rendering and geometry reconstruction. The reconstruction is fully unsupervised. We conduct extensive experiments on data from DTU and ShapeNet datasets, in which the method decomposes scenes into reasonable parts, outperforming existing state-of-the-art approaches.
Index Terms:
Part-aware reconstruction, Hybrid representation, 2D Gaussians, Superquadrics.
{strip}[Uncaptioned image]
Figure 1: We propose a pipeline to reconstruct the scene in a part-aware manner. Prior works [1, 2] model the scene using primitives, while the proposed method can further model precise geometry details and textures. Here, EMS [1] takes point cloud inputs and reconstructs non-textured primitives.
I Introduction
3D reconstruction from multi-view images is a long-studying problem in 3D vision and computer graphics. Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes, which is different from human perceptions. Humans understand 3D scenes or objects as different semantic parts [3]. For example, when we see a scene, we are building high-level structural information such as scene graphs in our minds, not point clouds or detailed geometry. Thus, we propose to learn part-aware scene reconstruction to decompose the scenes into different individual semantic parts or shapes, which would be more helpful for tasks including scene manipulation/editing, scene graph generation, etc.
There are prior methods [4, 5, 6, 7, 8, 9, 1, 10, 11, 12, 13] working on this topic, while they mainly learn the goal with 3D supervisions and cannot preserve accurate geometries, which leaves inconvenience to apply in realistic scenarios. While neural radiance field (NeRF) [14] has demonstrated the potential of reconstructing textured 3D scenes from multi-view images, some approaches [15, 16] try to learn part-aware objects via NeRFs. For example, Part-NeRF [15] represents an object as multiple neural radiance fields. However, the complex composition of multiple neural fields increases learning difficulty, resulting in subpar rendering outcomes and structurally insignificant information. Another state-of-the-art approach DBW [2] decomposes the scenes into multiple blocks represented by superquadric primitives [6], optimizing their parameters and UV texture maps to minimize rendering loss. Although they succeed in decomposing 3D scenes into reasonable parts, learning accurate detailed geometry and texture maps still proves challenging, or not included in the pipelines. Segmenting the scene into reasonable parts while preserving high-quality geometry details and textures is essential for good geometry reconstruction and rendering quality.
To tackle the problem, we propose integrating superquadrics and 2D Gaussians [17] to reconstruct part-aware and well-textured scenes, culminating in a hybrid representation. The hybrid representation retains the primitive property of superquadrics, where we assume each part should be a basic shape, and shapes that cannot be fitted as a superquadric should be segmented into multiple parts. On the other hand, Gaussian splatting, renowned for its superior rendering capabilities in recent studies [18, 19, 20] and efficient training, is incorporated to capture intricate texture details, speeding up the reconstruction and improving the rendering quality significantly.
Specifically, in the hybrid representation, Gaussian ellipses are attached to the surface of each superquadric mesh. We adopt 2D Gaussians and share their pose parameters with the attached superquadric mesh. Compared to the vanilla Gaussian [18] that fills up the whole occupancy with individual-parameter 3D Gaussians, our method is more compact and efficient. In this way, the parameters to optimize include global parameters controlling the shape, pose, and opacity of the hybrid block, as well as local parameters of spherical harmonic coefficients of the Gaussian ellipses. In the training stage, the parameters are optimized mainly by re-rendering loss without additional supervision. We observe that image rendering loss tends to settle into local minima during the deformation of superquadrics. To address this, we design several regularizers to enforce global consistency between the 3D representation and the input 2D information. This allows us to segment parts in a fully unsupervised manner, achieving block-level reconstruction.
At last, to model the accurate geometry, we conduct a point-level refinement, freeing the 2D Gaussians to get out of the surface, thereby improving the modeling of irregular shapes.
We summarize the contributions as follows:
  • A novel hybrid representation is introduced to learn part-aware 3D scenes, where primitives model the parts and Gaussians model the textures. The hybrid representation combines the benefits of both superquadrics and Gaussian splitting, enabling accurate part reconstruction and efficient high-quality rendering.
  • An end-to-end unsupervised pipeline is proposed for block-level and point-level part-aware reconstruction, with novel regularizers to optimize the superquadric meshes and 2D Gaussians simultaneously.
  • Extensive experiments on public datasets demonstrate the superiority of the proposed method in part-aware reconstruction, working for part segmentation of objects and scenes, outperforming several state-of-the-art methods.
Figure 2: Overview of the proposed method. It takes multi-view images as input to learn a parametric hybrid representation of superquadrics and 2D Gaussians. The representation initializes from random superquadrics and is gradually optimized during training to get a block-level reconstruction. At last, we free the constraints of Gaussians to model the detailed geometry to get a point-level reconstruction.
II Related Works
In this section, we first introduce classic multi-view reconstruction techniques. Next, we provide an overview of the advancements in primitive-based 3D representation. Additionally, we discuss related works on surface-aligned Gaussian splatting, which inspired our hybrid representation.
II-A Multi-View Reconstruction
Before the era of deep learning, multi-view stereo-based methods [21, 22, 23, 24] adopt a standard pipeline that consisting of image feature matching, depth prediction, and fusion from different viewpoints. Recent neural approaches [25, 26, 27] combining implicit representation and neural rendering outperforms previous methods in high-fidelity reconstruction. These techniques typically represent surfaces implicitly via a multi-layer perceptron (MLP) and extract surfaces post-training using the Marching Cubes algorithm [28]. Additionally, some works [29, 30] use volumetric representations to achieve fast learning. 3D Gaussian Splatting (3D-GS) [18] was introduced as a new technique integrating point-based rendering and splitting. Due to its rendering speeds and quality never seen before, Gaussian splatting has sparked a research boom and has been applied in 3D reconstruction [17, 31, 32]. Our work can be seen as an end-to-end approach to part-aware multi-view reconstruction, whose goal is to output the 3D reconstruction of parts from calibrated multi-view images.
II-B Primitive-based 3D Representation
Learning structured shape representations aims to decompose objects or scenes into semantically coherent geometric primitives, providing a crucial pathway to shape understanding and generation [33, 34, 35, 36, 37, 38]. The earliest classical approaches, dating back to the last century, such as Blocks World [39] and Generalized Cylinders [40] emphasize compactness as the central goal. However, these ideas struggle to handle the complexity of real-world data.
A substantial amount of prior work has focused on processing 3D data inputs, such as point clouds, meshes, and voxel grids. These methods decompose the 3D data into various types of primitive ensembles, including cuboids [4, 41], superquadrics [42, 43, 9, 1], and convex shapes [44, 45]. Recent advancements include MonteBoxFinder [41], which utilizes a RANSAC-based [46] approach with cuboids to represent the input point cloud, combined with Monte Carlo Tree Search for optimal selection. Alternatively, EMS (Expectation, Maximization, and Switching) [1] employs superquadrics to represent geometric primitives and proposes a probability-based approach to recover these primitives from the point cloud. However, such methods often rely on perfect 3D data, rarely available in real-world scenarios. In practice, 3D data is frequently incomplete or noisy, especially when reconstructed from 2D images. This noise and missing data significantly hinder the practical application of these methods.
There are only a few studies [15, 2, 16, 47] that attempted to create structure-aware 3D representations directly from images without relying on 3D supervision. For example, PartNeRF [15] uses ellipses to achieve shape abstraction, modeling each part separately using a neural radiation field. However, using ellipses makes PartNeRF more likely to decompose shapes based solely on spatial locations, which limits its ability to recover meaningful structural information. ISCO [47] and DBW [2] propose decomposing shapes using 3D superquadrics, which allows for meaningful shape decomposition. Nonetheless, the simplicity of their shape parameters struggles to capture complex geometries, leading to poor reconstruction accuracy and challenges in rendering high-quality novel view synthesis. Recently, DPA-Net [16] has made great progress on 3D shape abstraction from sparse views. However, this approach generates redundant parts and struggles with realistic texture rendering. Our approach introduces a hybrid representation to learn part-aware 3D scenes while maintaining accurate structural reconstruction and efficient, high-quality rendering capabilities.
II-C Surface-Aligned Gaussian Splatting
Gaussian splatting [18] has been rapidly adopted in multiple fields because of its remarkable capability in rendering. However, the challenge of unconstrained Gaussians resulting in inaccurate geometry during reconstruction has prompted several studies [48, 49, 50, 51, 17] to align Gaussians with scene surfaces. Among them, SuGaR [48] addresses this by using flat 3D Gaussians to enforce the alignment with the scene surface during optimization, minimizing the difference between the signed distance functions (SDF) of the desired Gaussians and the actual Gaussians. GaMeS [51] introduces a hybrid representation of Gaussians and mesh, where Gaussians are attached to the triangular facets of the mesh. It allows for mesh-like adaptation and modification of Gaussian components. Similarly, Gao et al. [50] propose a mesh-based Gaussian Splatting to achieve large-scale deformation effects on objects. NeuSG [49] combines 3DGS with implicit SDF networks, optimizing the Gaussian components and the network parameters jointly. Recently, 2DGS [17] proposes using 2D Gaussians instead of 3D Gaussians, leveraging 2D surface modeling to significantly enhance the geometric quality of scenes. Inspired by them, our work proposes a representation that combines 2D Gaussians with superquadrics, ensuring that the Gaussians conform to the surface of the superquadrics.
III Method
Given a set of calibrated multi-view images, we aim to learn the part-aware 3D geometry and texture information based on Gaussian splatting. We observe that a part-aware scene can intuitively be regarded as a composition of multiple basic shapes (geometric primitives). Therefore, we design a primitive-based hybrid representation to describe the 3D scene, as introduced in Section III-A. In Section III-B, this hybrid representation can be optimized to fit primitives as different parts, as stage 1. In stage 2, we refine the positions of 2D Gaussians to model the precise geometry of the original shape.
III-A Hybrid Representation
As shown in Fig. 2, the proposed hybrid representation combines superquadrics with 2D Gaussian ellipses in a compact form. Superquadrics are efficient for modeling 3D scenes due to their primitive properties, but they fall short in capturing detailed geometry and texture information. Conversely, the 3D/2D Gaussian representation [18, 17] excels in photorealistic rendering but is cumbersome and lacks structural awareness.
To leverage the strengths of both representations, we attach 2D Gaussian ellipses to the surface of the superquadric mesh. This approach retains the superquadric’s advantage of effectively parsing a 3D scene into distinct parts. At the same time, the Spherical Harmonics of 2D Gaussians enable the rendering of complex textures by Gaussian splatting, addressing the limitations in texture learning. By sharing the pose parameters between superquadrics and 2D Gaussians, we achieve a more efficient representation.
III-A1 The Parametric Model
The proposed hybrid model is parametric, controlled by primitive parameters and Gaussian parameters. Considering a 3D scene $\mathcal{S}$, we parse it into multiple superquadrics, i.e. blocks, in the first stage. Each block is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$. Here, $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks. The hybrid blocks are initialized using manually designed parameters that control pose, opacity, scale, shape, and texture. These parameters are optimized through a multi-view rendering loss to learn the scene.
Shape and Scale Parameters. For each hybrid block $\mathcal{B}_{i}$, its shape is defined by the superquadric mesh. As introduced in [6], a superquadric comprises a series of parametric surfaces controlled by several continuous parameters. More concretely, there are 2 parameters $\epsilon_{1},\epsilon_{2}$ defining the superquadric mesh and 3 anisotropy parameters $s_{1},s_{2},s_{3}$ scaling the 3D axes. Similar to the position of each vertex of an icosphere, we can acquire the 3D position of the superquadric vertex by:
$\scriptstyle\mathbf{v}=[s_{1}\cos^{\epsilon_{1}}(\theta)\cos^{\epsilon_{2}}(\varphi);s_{2}\sin^{\epsilon_{1}}(\theta);s_{3}\cos^{\epsilon_{1}}(\theta)\sin^{\epsilon_{2}}(\varphi)],$ (1)
where $\theta$ and $\varphi$ are the azimuth and elevation defined in the spherical coordinate system. The shape and scale parameters are optimized during the training stage, thus controlling the block deformation to learn part-aware information.
Pose Parameters. The initial hybrid blocks are randomly sampled and disorganized in the space, so it is necessary to align their poses to fit within the 3D scene. The vertex position acquired in Eq. 1 is in the local coordinate system of the hybrid block, so we need to use the pose parameters to transform the position into the world space. We define the pose parameter of the $i$-th hybrid block as the combination of rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$. Hence, the vertex position in the world space is calculated by: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $i,j$ denotes the indexes of block and vertex respectively. In general, the Gaussians placed on the superquadric surface also have their individual poses, which causes a large amount of parameters. To address this issue, we propose to share the pose of the superquadric with the Gaussians. Given a posed superquadric, we directly determine the center points of the Gaussians through uniform sampling on the triangular faces and let the normal of 2D Gaussians be the same as the faces. Inspired by GaMeS [51], we assign the rotation matrix $\mathrm{R}_{v}$ and scaling $\mathrm{S}_{v}$ of Gaussians based on the vertices’ positions. For a given triangular face $V$ and its three vertices $v_{1}$, $v_{2}$, $v_{3}\in\mathbb{R}^{3}$, we define the orthonormal vectors $r_{1}$, $r_{2}$, $r_{3}\in\mathbb{R}^{3}$ to construct the rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$. The first vector $r_{1}$ is aligned with the normal vector of $V$ to ensure the Gaussians align with the triangular surface. The vector $r_{2}$ is defined as the vector from the center $m=\text{mean}(v_{1},v_{2},v_{3})$ of the triangle to the first vertex $v_{1}$:
$r_{2}=\frac{v_{1}-m}{||v_{1}-m||}.$ (2)
The last rotation vector $r_{3}$ is obtained by orthonormalizing the vector from the center to the second vertex concerning the existing $r_{1}$ and $r_{2}$ :
$r_{3}=\frac{{ort}(v_{2}-m;r_{1},r_{2})}{||{ort}(v_{2}-m;r_{1},r_{2}||},$ (3)
where ${ort}$ represents the orthonormalization step in the Gram-Schmidt [52]. For the scaling parameters of the 2D Gaussians, we use:
$\mathrm{S}_{v}=\text{diag}(s_{v}^{2},s_{v}^{3}),$ (4)
where $s_{v}^{2}=c||m-v_{1}||$, and $s_{v}^{3}=c<v_{2},r_{3}>$, and $c$ is a size control hyperparameter. We place a fixed number of $n$ Gaussians on each triangular face.
This design avoids the additional cost of optimizing the pose and covariance matrix of Gaussians, making it more efficient for representing the scene. The pose parameters are optimized via gradient descent so that the hybrid blocks can fit the scene.
Opacity Parameters. We define the total number of hybrid blocks as $M$. However, in practice, a typical scene does not contain exactly $M$ blocks, so we retain only the meaningful ones. Therefore, we utilize a learnable parameter $\tau_{i}$ to represent each block’s opacity. In the optimization process, only blocks with $\tau_{i}$ greater than a certain threshold are retained. Note that Gaussians on the same block share the same $\tau$ for opacity in rasterization.
Texture Parameters. The texture information is modeled using 2D Gaussians positioned on the surface of the superquadric mesh. Spherical harmonics control the texture of each Gaussian ellipse. Therefore, we define the coefficients of the spherical harmonics as texture parameters and optimize them for rendering view-dependent images.
III-A2 Rendering
With Gaussians attached to the surface of each block, we achieve view-dependent rendering through tile-based rasterization as Gaussian Splatting [18, 17]. Given a view, Gaussians on the superquadric surface are projected onto the image space, forming an RGB image. Initially, the screen space determines the bounding box for each Gaussian. Subsequently, these Gaussian ellipses are sorted according to their depths of center to the image plane. Finally, volumetric alpha compositing [53] is utilized to integrate the alpha-weighted RGB values for each pixel.
To formulate the process, considering a viewing transformation ${W}$, the covariance matrix ${\Sigma}_{i}^{{}^{\prime}}$ of $i$-th Gaussian in the camera coordinate system is calculated by:
${\Sigma}_{i}^{{}^{\prime}}={J}{W}{\Sigma}{W}^{T}{J}^{T},$ (5)
where ${J}$ is the Jacobian of the affine approximation of the projective transformation, and ${\Sigma}$ is the covariance matrix of the Gaussian ellipse. Note that the last row and column of ${\Sigma}$ are omitted since we adopt 2D Gaussians. Following alpha compositing, we first calculate an alpha value for each Gaussian ellipse:
$\alpha_{i}=\tau_{i}\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{u}_{i})^{T}{\Sigma_{i}^{{}^{\prime}}}_{-1}(\mathbf{x}-\mathbf{u}_{i}))).$ (6)
Here, $\mathbf{u}_{i}$ is the center coordinate of the projected Gaussian ellipse, and $\tau_{i}$ is the opacity of the block where the $i$-th Gaussian is located. The calculated alphas are sorted according to their depths from the image plane. Meanwhile, we can acquire the color value $c_{i}$ from the spherical harmonics of $\mathcal{N}$ ordered points, thus obtaining the rendering RGB value:
$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (7)
This process is differentiable and can optimize the hybrid representation through gradient descent, and the optimization/learning is introduced in Sec. III-B.
III-B Optimizing the Hybrid Representation
In this section, we introduce how to optimize the hybrid representation. By minimizing the rendering loss across multi-view images only, we found it unstable to position the hybrid blocks (primitives) accurately. Thus, we introduce several regularizations and perform two optimization stages (block-level and point-level optimizations).
III-B1 Block-level Optimization
Besides the image rendering loss, we employ regularizers consisting of coverage, overlap, parsimony, and opacity entropy. The coverage loss encourages hybrid blocks to cover only the meaningful regions; the overlap loss prevents different blocks from overlapping; the parsimony loss regularizes the number of existing blocks; the opacity entropy encourages opacities of hybrid blocks close to binary.
Rendering Loss. Based on Eq. 7, we can obtain the rendered image $\hat{\mathbf{I}}$ associated with a given camera pose. We calculate the rendering loss between $\hat{\mathbf{I}}$ and the ground truth image $\mathbf{I}$ by associating $L_{1}$ with a D-SSIM term[18]:
$L_{\text{ren}}=(1-\lambda)\ L_{1}+\lambda L_{\text{D-SSIM}}.$ (8)
Coverage Loss. This loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending outside the object. To determine the occupancy relationship in 3D space, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=1-\Psi_{i}(x)$. Here, $\Psi_{i}$ denotes the superquadric inside-outside function [6] associated to the block $i$. Consequently, $D_{i}(x)\geq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)<0$ if $x$ is outside of the block. Similar to NeRF [14], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses. Given the object mask, we associate each ray $r\in\mathcal{R}$ with a binary label, such that $l_{r}=1$ if the ray $r$ is inside the mask, otherwise $l_{r}=0$. The coverage loss is defined as:
${\mathcal{L}}_{\text{cov}}({\mathcal{R}})=\sum_{r\in{\mathcal{R}}}l_{r}L_{\text{cross}}(r)+(1-l_{r})L_{\text{non-cross}}(r).$ (9)
Here, $L_{\text{cross}}$ encourages rays to intersect with blocks , while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks. Intuitively, this means that there is at least one in sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not lie within any block. This process can be expressed as:
$L_{\text{cross}}(r)=\text{ReLU}\Big{(}\min_{i\in\mathcal{M}}\min_{x_{j}^{r}\in{\mathcal{X}}_{r}}D_{j}(x_{j}^{r})\Big{)},$ (10)
$L_{\text{non-cross}}(r)=\text{ReLU}\Big{(}\max_{i\in\mathcal{M}}\max_{x_{j}^{r}\in{\mathcal{X}}_{r}}-D_{i}(x_{j}^{r})\Big{)},$ (11)
where $x_{j}^{r}$ means the $j$-th sampled points along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$.
Overlap Loss. We add a regularization term to encourage individual blocks to be non-overlapping. Considering the difficulty in directly calculating the overlap of locks, we use a Monte Carlo method similar to  [2]. Practically, we sample multiple 3D points in space and penalize each point for being inside more than $k$ blocks. Based on the definition of the superquadric inside-outside function, we devise a soft occupancy function:
$\mathcal{O}_{i}^{x}=\tau_{i}\text{ sigmoid}(\frac{1-\Psi_{i}(x)}{\gamma}),$ (12)
where $\gamma$ is a temperature parameter. Therefore, given a set of $N$ sampled 3D points $\Omega$, the overlap loss can be expressed as:
$L_{\text{over}}=\frac{1}{N}\sum_{x\in\Omega}\text{ReLU}\left(\sum_{i\in\mathcal{M}}\mathcal{O}_{i}^{x}-k\right).$ (13)
Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we add a loss term that regularizes the number of active blocks. This loss is defined as:
$L_{\text{par}}=\frac{1}{M}\sum_{i\in\mathcal{M}}\sqrt{\tau_{i}}.$ (14)
Opacity Entropy Loss. Naturally, during optimization, the opacity of blocks inside the object region tends to approach 1, while the opacity of blocks outside the object region tends to be 0. To facilitate this, we associate the opacities of the blocks with the labeled rays. Specifically, we employ a cross-entropy loss, $L_{ce}$, between the opacity and mask labels as follows:
${\mathcal{L}}_{\text{opa}}({\mathcal{R}})$ $=\frac{1}{|{\mathcal{R}}|}\sum_{r\in{\mathcal{R}}}L_{ce}\left(\max_{i\in\mathcal{M}}\tau_{i}(x^{r}),l_{r}\right)~{},$ (15)
$\left\{x^{r}\in{\mathcal{X}}_{r}\mid D_{i}(x^{r})\leq 0\right\}.$
Thus, the total loss is the weighted summation of the above terms:
$L=L_{\text{ren}}+\lambda_{\text{cov}}L_{\text{cov}}+\lambda_{\text{over}}L_{\text{over}}+\lambda_{\text{par}}L_{\text{par}}+\lambda_{\text{opa}}L_{\text{opa}}.$ (16)
Adaptive Number of Blocks. Given that the number of parts varies by scene, we allow the number of blocks to adjust during optimization adaptively. When the opacity of an existing block falls below a threshold $t$, this block is immediately removed. Additionally, we explore a block-adding mechanism using point clouds that are randomly generated within the scene area. Specifically, we employ DBSCAN [54] to cluster point clouds within the perspective of the object masks and not covered by any blocks. A new primitive is then introduced at the center of each point cloud cluster. The other parameters of these new blocks are initialized randomly.
At last, block-level reconstruction gets a 3D scene composited by several superquadrics, as shown in Fig. 2.
III-B2 Point-level Optimization
The primitive-based hybrid representation effectively decomposes the shape into parts but underperforms with complex objects. To address this, we perform a refinement stage for precise geometry fitting. In this stage, we free the constraint that Gaussian ellipses must be at the face of the superquadric mesh, allowing Gaussians to move freely to fill in the complex regions. Moreover, to prevent Gaussian ellipses from one block passing through another, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
$L_{\text{enter}}=\frac{1}{N}\sum_{x\in\Omega}\sum_{m\in\mathcal{M}\setminus\{\delta\}}\text{ReLU}(-D_{m}(x)),$ (17)
where $\Omega$ is the sampled Gaussian set, and $\delta$ represents the identifier of the block to which the Gaussian $x$ belongs. After this stage, as shown in Fig. 2, the reconstruction is more precise and better fits the original shape.
Figure 3: Qualitative comparisons on DTU [55]. We compare our approach with state-of-the-art baselines on the DTU dataset with the background removed. The last two columns show our block-level and point-level reconstructions, respectively. Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
Figure 4: Qualitative comparisons on ShapeNet [55]. We compared our approach with state-of-the-art baselines across four categories. The last two columns display our block-level and point-level reconstructions, respectively. Our method uniquely provides reasonable 3D part decomposition while simultaneously capturing detailed geometry.
Figure 5: Qualitative results on BlenderMVS [56] and self-capatured data. We demonstrate the RGB renderings and decomposed parts from novel views across a variety of objects. The first five examples are from the BlenderMVS dataset, and the remaining examples are from our own captured scenes.
Figure 6: Ablation studies on key strategies. The block-level visual comparisons illustrate the impact of adopting our proposed strategy. The first row indicates results without the strategy, while the second row shows results with the strategy implemented.
TABLE I: Quantitative comparison on DTU [55]. The Chamfer Distance (CD) between the 3D reconstruction and the ground-truth is reported in 15 scenes. The best results are bolded, and the average numbers of primitives found (#P) that are greater than 10 are underlined.
Chamfer Distance (CD) per scene Mean Mean
Method Input S24 S37 S40 S55 S63 S65 S69 S83 S97 S105 S106 S110 S114 S118 S122 CD #P
EMS [1] GT point cloud 6.32 3.54 2.99 4.30 4.16 4.01 3.75 3.24 4.97 4.34 4.16 7.62 7.58 4.46 4.03 4.65 7.7
MBF [41] GT point cloud 3.12 2.66 3.84 2.54 1.59 2.11 2.19 2.01 2.32 2.45 2.17 2.12 3.83 2.02 2.55 2.50 34.1
EMS [1] + Neus [27] Iamge 5.99 5.56 4.43 4.32 5.42 6.14 3.75 3.96 4.63 4.34 5.88 5.11 4.29 4.83 3.53 4.97 8.87
MBF [41] + Neus [27] Image 2.69 3.37 3.22 2.69 3.63 2.60 2.59 3.13 2.85 2.51 2.45 3.72 2.24 2.49 2.52 2.85 46.7
PartNeRF [15] Image 9.38 10.46 9.08 8.63 6.04 7.25 7.22 9.15 8.72 10.01 6.72 9.85 7.85 8.68 9.21 8.54 8.0
DBW [2] Image 5.41 8.35 1.57 3.08 3.40 4.15 7.46 3.94 6.63 4.85 4.38 4.65 6.29 4.34 3.04 4.76 4.8
Ours (Block-level) Image 5.68 4.91 1.85 2.61 3.75 4.66 3.75 7.57 4.27 4.38 3.49 4.48 3.61 4.21 3.70 4.19 5.9
Ours (Point-level) Image 0.95 1.38 0.55 0.71 1.02 1.39 1.16 1.33 1.49 0.80 0.95 1.31 0.69 1.20 0.82 1.05 5.9
TABLE II: Quantitative comparisons on DTU [55]. We compare our method to PartNeRF, DBW, and 2DGS in terms of PSNR, SSIM, LIPPS, Chamfer Distance, and reconstruction time. All running time measurements were conducted on a single GTX RTX3090 GPU. Excluding the results of 2DGS, the best results are in bold font, and the second-best results are underlined.
Method Structure CD $\downarrow$ PSNR $\uparrow$ SSIM  $\uparrow$ LIPPS $\downarrow$ Time $\downarrow$
2DGS [17] 0.96 34.07 0.99 0.019 $\sim$ 18 mins
PartNeRF [15] 9.59 17.97 0.77 0.246 $\sim$ 8 hours
DBW [2] 4.73 16.44 0.75 0.201 $\sim$ 2 hours
Ours (Block-level) 4.19 19.84 0.82 0.189 $\sim$ 30 mins
Ours (Point-level) 1.05 35.04 0.99 0.015 $\sim$ 50 mins
TABLE III: Quantitative comparison on ShapeNet [57]. The Chamfer Distance (CD) between the 3D reconstruction and the ground-truth is reported on objects of four categories. The best results are bolded, and the second-best results are underlined.
Chamfer Distance  $\downarrow$ Primitives (#P ) Mean Mean
Method Input Airplane Table Chair Gun Airplane Table Chair Gun CD #P
EMS [1] GT point cloud 3.40 6.92 19.0 2.02 9.4 7.88 10.3 8.4 7.84 9.0
PartNeRF [15] Image 2.29 2.77 2.30 2.46 8.0 8.0 8.0 8.0 2.46 8.0
DBW [2] Image 3.61 7.33 6.19 2.09 2.7 5.2 3.6 3.3 4.81 3.7
Ours (Block-level) Image 2.47 2.15 2.32 1.78 3.9 6.6 7.6 5.0 2.18 5.8
Ours (Point-level) Image 1.29 1.72 0.94 1.07 3.9 6.6 7.6 5.0 1.25 5.8
TABLE IV: Quantitative comparison on ShapeNet [57]. We report the PSNR, SSIM, LIPPS, Chamfer Distance, and reconstruction time for our method, PartNeRF, as well as DBW and 2DGS across four categories. Excluding the results of 2dgs, the best results are bolded, and the second-best results are underlined.
Chamfer Distance  $\downarrow$ PSNR  $\uparrow$ SSIM  $\uparrow$ LIPPS  $\downarrow$
Method Structure Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun
2DGS [17] 1.47 2.37 0.50 1.03 40.89 39.80 39.05 41.74 0.994 0.990 0.990 0.994 0.009 0.027 0.017 0.009
PartNeRF [15] 2.29 2.77 2.30 2.46 19.63 20.66 19.08 21.97 0.898 0.855 0.875 0.916 0.086 0.161 0.136 0.083
DBW [2] 3.61 7.33 6.19 2.09 26.11 23.84 20.25 28.72 0.950 0.915 0.892 0.960 0.074 0.136 0.132 0.042
Ours (Block-level) 2.47 2.15 2.32 1.78 26.93 26.92 23.91 27.97 0.936 0.911 0.891 0.942 0.082 0.139 0.126 0.072
Ours (Point-level) 1.29 1.72 0.94 1.07 41.18 36.80 36.07 39.51 0.992 0.973 0.977 0.989 0.014 0.070 0.038 0.021
IV Experiments
IV-A Experimental Settings
Datasets. We conduct evaluation experiments on two public datasets: DTU [55] and ShapeNet [57]. DTU is an MVS dataset comprising 80 forward-facing scenes, with 49-60 images captured for each scene. For evaluation, we utilize 15 scenes used in 2DGS [17]. We downsample the images to a resolution of 400 $\times$ 300 for efficiency. To validate the effectiveness of our method in segmenting parts of man-made objects, we create a ShapeNet dataset that includes four categories: Chair, Table, Gun, and Airplane. Each category comprises 15 distinct objects, offering diverse instances for comprehensive evaluation. We randomly sample 100 camera poses on the sphere centered on each shape and render images at a resolution of 400$\times$400. The rendered images are equally split into training and testing sets. Additionally, to explore our method’s potential in handling real-life data, we present qualitative results on the BlenderMVS dataset [56] and self-captured scenes. For BlenderMVS data, we utilize standard camera poses. For self-captured scenes, we recover the camera poses using COLMAP [21] and normalize them following IDR [58]. All images are resized to 400 $\times$ 300 pixels. Object masks are obtained using Grounded SAM [59].
Baselines. We compare our method with four state-of-the-art works on 3D shape decomposition: EMS [1], MonteboxFinder (MBF) [41], PartNeRF [15] and DBW [2]. MBF and EMS are applied to point clouds, utilizing cuboids and superquadrics, respectively, to fit the 3D points. The input point clouds are sampled from either the original 3D ground truth (GT) or the mesh reconstructed by the state-of-the-art MVS method Neus [27]. We adhere to the testing procedure outlined by DBW [2] and randomly sample 5K and 200K points from the GT meshes as the input for EMS and MBF, respectively. Similar to our problem setting, PartNeRF and DBW use images directly as input to construct structural 3D representations. To ensure a fair comparison, we discarded the pre-defined ground plane in DBW since the test scenario consists solely of foreground objects. For PartNeRF, we perform instance-specific training and set the same number of parts as in our method. Additionally, we compare our method with 2DGS [17], the state-of-the-art Gaussian Splatting reconstruction method.
Metrics. We evaluate from 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time. The 3D reconstruction quality is measured by the Chamfer Distance (CD) between the recovered geometry and GT, reflecting the accuracy of 3D reconstructions. View synthesis uses three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [60], and Learned Perceptual Image Patch Similarity (LPIPS) [61]. Shape parsimony and reconstruction time are measured by the number of parts and running time, where fewer parts facilitate easier part editing or manipulation.
IV-B Implementation Details
We improve the 2DGS [17] architecture by adding a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq. 17 during refinement optimization. We employ regularization in 2DGS to achieve better geometric reconstruction, including depth distortion maps, depth maps, and normal maps. Additionally, we introduce a mask cross-entropy to filter out extra noise Gaussians. To extract the meshes from 2D Gaussians, we use truncated signed distance fusion (TSDF) to fuse the rendered depth map, utilizing Open3D [62]. In the ShapeNet benchmark, we observed a common degradation in the mesh reconstruction by 2DGS, where the meshes of objects contain many holes. This issue arises due to the lack of textures in ShapeNet images, causing 2DGS to use large-scale Gaussians. This leads to depth inconsistencies from multiview, resulting in holes in the constructed mesh. To address this, we apply scale regularization to limit the scale of large Gaussian components. We validate this design in Sec. IV-D.
The same hyperparameters are used for all experiments. We set the initial number of primitives $M$ to 8. In the hybrid representation, each superquadric mesh is a level-2 icosphere (320 triangular face). Each triangular face contains 100 Gaussians, with a scale parameter $c$ of 0.15 and a block filter threshold $t$ of 0.2. The temperature parameter $\gamma$ in Eq. 12 is 0.005 and the overlapping number $k$ in Eq. 13 is 1.95. We train the hybrid representation for 30k iterations, followed by refinement optimization for another 30k iterations. During block-level optimization, the block-adding operation is carried out at the 5000th and 10000th iterations. The point clouds used for block-adding are randomly initialized within the space $[-1,1]^{3}$. The number of sampled points in each ray is 2048 in Eq. 9 and Eq. 15. The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.01, and 0.01, respectively.
IV-C Evaluations
IV-C1 DTU Benchmark
We utilize the official DTU Chamfer Distance (CD) evaluation [55] to compare the performance between our method and these state-of-the-art 3D decomposition techniques, as shown in Tab. I. Additionally, we present the input type and the average number of primitives used. Our method consistently outperforms in all scenes compared to prior works. Even without refinement optimization, our method still provides satisfactory CD errors. PartNeRF has the largest CD errors in the DTU dataset. We speculate that this is due to PartNeRF’s difficulty in handling forward scenes. In contrast, MBF achieves a low CD error but at the cost of using significantly more primitives, leading to over-decomposition and the inclusion of meaningless parts. Although EMS and DBW utilize a reasonable number of primitives, they yield higher CD errors than the proposed method.
We compare the quality of view synthesis and reconstruction time with PartNeRF, DBW, and 2DGS in Table II. Even without point-level optimization, our method achieves superior image rendering quality compared to other state-of-the-art methods. With the addition of refinement, both image rendering quality and reconstruction accuracy are further improved, nearing the performance of the state-of-the-art Gaussian-based reconstruction method, 2DGS [17]. However, 2DGS can only produce unstructured meshes. Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than the implicit reconstruction method PartNeRF and over 2X faster than DBW.
Figure 3 presents visual comparisons of our method with existing state-of-the-art baselines. Even with ground-truth 3D points, EMS and MBF still struggle to extract reasonable compositions. PartNeRF reconstructs reasonable geometry but fails to decompose the scenes meaningful. DBW experiences frequent failures, particularly when parts of the geometry are not reconstructed due to the absence of primitives in these regions at initialization. In contrast, our approach consistently produces interpretable 3D decompositions (second-to-last column) and, through refinement, achieves more detailed structures (last column).
IV-C2 ShapeNet Benchmark
We evaluate our approach on four categories of ShapeNet [57]. Since MBF [41] fails on nearly all objects of the ShapeNet benchmark, we compare only against EMS [1], PartNeRF [15] and DBW [2]. Fig. 4 presents the qualitative comparisons of the 3D decomposition of eight object shapes, with two instances per category. Tab. III and Tab. IV deliver detailed quantitative comparisons on reconstruction quality, novel view synthesis, and primitives parsimony. For clarity, the chamfer distance is multiplied 100 times in the Shapenet benchmark.
EMS yields large Chamfer distances, especially in the chair and table categories, despite fitting superquadrics from GT 3D points. Similarly, DBW often fails to recover complete geometry in these categories, resulting in significant reconstruction errors. However, it performs better with guns and airplanes due to their simpler shapes. The non-convex nature of chairs and tables poses substantial challenges for EMS and DBW in generating reasonable geometry. PartNeRF benefits from NeRF’s geometric reconstruction capabilities to produce accurate geometry and good novel view renderings. However, it has two main disadvantages. On the one hand, PartNeRF cannot adaptively adjust the primitives, resulting in a fixed number of parts. Estimating the number of parts in advance is impractical. On the other hand, its use of ellipsoidal primitives limits its ability to decompose meaningful parts, such as separating the chair seat and tabletop into several components, as shown in the three columns of Fig. 4.
Conversely, the superior performance of our approach is demonstrated across all categories, visually and quantitatively. In addition, thanks to Gaussian’s powerful rendering capabilities, clear advantages in novel view synthesis are observed over PartNeRF and DBW, as shown in Tab. IV. The quantitative comparisons against 2DGS [17] further validate the capability of our method on reconstruction accuracy. Our method outperforms 2DGS in the chair and airplane categories and achieves comparable performance in the other two categories.
TABLE V: Ablation study on the ShapeNet [57] chair benchmark. We report Chamfer Distance (CD), rendering metrics and the number of primitives (#P) of block-level reconstruction and point-level reconstruction. The best results are in bold font, and the second-best results are underlined.
Block-level Point-level
[…]
Complete model $2.32$ $23.91$ $0.891$ $0.126$ $\underline{0.94}$ $\underline{36.07}$ $0.977$ $0.038$ $6.6$
w/o $L_{\text{over}}$ $\underline{2.11}$ $\underline{24.13}$ $\underline{0.896}$ $\underline{0.122}$ $1.34$ $36.04$ $\underline{0.977}$ $\underline{0.037}$ $7.6$
w/o $L_{\text{opa}}$ $5.72$ $20.12$ $0.853$ $0.174$ $1.06$ $35.60$ $0.975$ $0.040$ $3.8$
w/o $L_{\text{cov}}$ $3.21$ $22.96$ $0.896$ $0.134$ $1.34$ $36.00$ $0.976$ $0.038$ $8.53$
w/o Adaptive $3.16$ $22.74$ $0.880$ $0.141$ $1.05$ $35.74$ $0.974$ $0.040$ $6.7$
w/o $L_{\text{par}}$ 1.91 24.47 0.898 0.119 0.93 36.18 0.977 0.037 $10.1$
IV-C3 Real-life Data
To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from BlenderMVS dataset [56] and a self-captured dataset. As shown in Fig. 5, our approach can robustly produce both realistic appearances and meaningful 3D decompositions across a variety of data types. Additionally, the distinct advantages of block-level and point-level reconstructions are clearly observable. For instance, in the sixth column, due to limited training views, point-level reconstructions based on 2DGS tend to generate artifacts, whereas block-level optimization produces accurate and smooth surfaces. Conversely, point-level reconstruction demonstrates superior capability in capturing intricate geometric details.
Figure 7: Impact of Gaussian scale regularization (SR). This figure demonstrates the effect of Gaussian scale regularization. As shown by the typically degraded mesh reconstruction from 2DGS [17], scale regularization effectively fixes the holes in the ShapeNet [57] dataset.
Figure 8: Impact of the initial number of primitives. We qualitatively demonstrate the impact of the initial number of primitives on the granularity of decomposition. A larger number of initial primitives results in a finer-grained decomposition, whereas a smaller number leads to a coarser decomposition.
Figure 9: Applications. Part-Aware Editing (top): After optimization, we can easily edit the scene by moving, adding, removing, or scaling specific parts. 3D Content Generation (bottom): By combining parts of different objects, we can effortlessly create new and meaningful 3D content.
IV-D Ablations
TABLE VI: Effect of the parsimony loss weight ($\lambda_{\text{par}}$) on DTU [55]. We report Chamfer Distance (CD), rendering metrics, and the number of primitives (#P) of block-level reconstruction and point-level reconstruction.
Block-level Point-level
[…]
$\lambda_{\text{par}}$ = 0.1 $6.46$ $16.10$ $0.721$ $0.255$ $1.08$ $34.30$ $0.987$ $0.018$ $2.7$
$\lambda_{\text{par}}$ = 0.01 (default) $4.19$ $19.84$ $0.820$ $0.189$ $1.05$ $35.04$ $0.988$ $0.015$ $5.9$
$\lambda_{\text{par}}$ = 0.001 $4.01$ $20.18$ $0.835$ $0.180$ $1.03$ $35.04$ $0.988$ $0.015$ $7.1$
TABLE VII: Effect of the initial number of primitives ($M$) on DTU [55]. We report Chamfer Distance (CD), rendering metrics, and the number of primitives (#P) of block-level reconstruction and point-level reconstruction.
Block-level Point-level
[…]
$M$= 4 $4.97$ $19.06$ $0.794$ $0.213$ $1.06$ $34.93$ $0.987$ $0.016$ $5.0$
$M$= 8 (default) $4.19$ $19.84$ $0.820$ $0.189$ $1.05$ $35.04$ $0.988$ $0.015$ $5.9$
$M$= 16 $3.99$ $20.87$ $0.83$ $0.176$ $1.07$ $35.05$ $0.988$ $0.015$ $8.3$
In this section, we first verify and analyze the critical strategies proposed by our algorithm: overlapping loss, opacity loss, coverage loss, adaptive number of primitives, and parsimony loss. By removing each strategy one at a time, we report the averaged quantitative performance of reconstruction on 15 instances of chair category in Tab. V and provide visual comparisons in Fig. 6.
As shown in Tab. V, VI, and VII, the point-level reconstruction accuracy and rendering quality do not differ significantly. This is because the refinement stage primarily assists in recovering the details of geometry rather than part decomposition. Therefore, in the subsequent analysis, we will focus on the block-level reconstruction results. Removing the overlapping loss improves reconstruction quality and reduces decomposition semantics while retaining more parts. Intuitively, more superquadrics will better fit a shape, but overlaps diminish the meaningfulness of decomposition, as shown in the first column of Fig. 6. Removing opacity loss drastically affects the number of primitives, reducing them to an average of only 3.8. Simultaneously, the CD increases by approximately twofold. This demonstrates the importance of opacity loss in controlling the existence of primitives and ultimately impacting reconstruction accuracy. In the visual example, the reconstructed chair is missing its legs and armrests. Coverage loss significantly involves reconstruction errors as the CD increases by 38%, ensuring the primitives stay in the correct position to fit the target best. As shown in the third column of Fig. 6, there is space in the chair seat that primitives do not cover without coverage loss. The adaptive number of primitives strategy fills in missing parts and improves the reconstruction metric by 36%. Omitting parsimony loss results in the lowest CD errors and best rendering metrics but produces 0.5 times more primitives than the full model, indicating redundancy and over-decomposition.
we also validate the effectiveness of Gaussian scale regularization in Fig. 7. We find that 2DGS [17] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF. The holes occur since large-scale Gaussians can’t support consistent depth rendering from different views. Scale regularization effectively fixes these holes by preventing the formation of large-scale Gaussians.
Lastly, we analyze the impact of two key hyperparameters of our method on decomposition granularity: the parsimony loss weight $\lambda_{\text{par}}$ and the initial number of primitives $M$. As shown in Tab.VI, $\lambda_{\text{par}}$ directly influences the number of primitives found (#P). Stronger regularization results in fewer primitives, while weaker regularization has the opposite effect. However, having too few primitives severely impacts reconstruction accuracy and rendering quality. In Tab.VII, it can be observed that as $M$ increases, both reconstruction and view synthesis performance slightly improve. Simultaneously, $M$ positively correlates with #P. This indicates that by adjusting the values of $M$ and $\lambda_{\text{par}}$, our approach can control the granularity of object or scene decomposition. Fig. 8 visually represents this impact.
IV-E Applications
Fig. 9 illustrates two applications of our method that original 3DGS and NeRF-based approaches do not support. Firstly, after optimization, we obtain the part decomposition, making it easy to edit specific parts of the object or scene, e.g., adding, moving, removing, or scaling sizes. Secondly, by combining parts of different objects, our method enables the creation of new high-quality 3D content. By replacing parts, we do not need to adjust the position and part connections manually. For man-made objects, the proposed method supports 3D shape generation within the same categories nicely.
V Conclusion
We introduce a hybrid representation of superquadrics and 2D Gaussians, to learn 3D scenes in a part-aware representation. Compared to prior works, the proposed method also keeps geometry details instead of shape abstractions, supporting high-level image rendering quality. The method gets state-of-the-art performance in comprehensive evaluations. One limitation is that the method takes background-free scenes as inputs, from segmentation tools such as SAM, as we care about the structure of the centered scene the most. In the future, we aim to explore how to model backgrounds and extend the work to address entire complex scenes.

### v2

@@
-Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics
+PartGS: Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics
 Abstract
 Low-level 3D representations, such as point clouds, meshes, NeRFs, and 3D Gaussians, are commonly used to represent 3D objects or scenes.
-However, humans usually perceive 3D objects or scenes at a higher level as a composition of parts or structures rather than points or voxels.
-Representing 3D as semantic parts can benefit further understanding and applications.
-We aim to solve part-aware 3D reconstruction, which parses objects or scenes into semantic parts.
-In this paper, we introduce a hybrid representation of superquadrics and 2D Gaussians, trying to dig 3D structural clues from multi-view image inputs.
+However, human perception typically understands 3D objects at a higher level as a composition of parts or structures rather than points or voxels.
+Representing 3D objects or scenes as semantic parts can benefit further understanding and applications.
+In this paper, we introduce PartGS, part-aware 3D reconstruction by a hybrid representation of 2D Gaussians and Superquadrics, which parses objects or scenes into semantic parts, digging 3D structural clues from multi-view image inputs.
 Accurate structured geometry reconstruction and high-quality rendering are achieved at the same time.
-We incorporate parametric superquadrics in mesh forms into 2D Gaussians by attaching Gaussian centers to faces in meshes.
-During the training, superquadrics parameters are iteratively optimized, and Gaussians are deformed accordingly, resulting in an efficient hybrid representation.
-On the one hand, this hybrid representation inherits the advantage of superquadrics to represent different shape primitives, supporting flexible part decomposition of scenes.
-On the other hand, 2D Gaussians are incorporated to model the complex texture and geometry details, ensuring high-quality rendering and geometry reconstruction.
-The reconstruction is fully unsupervised.
-We conduct extensive experiments on data from DTU and ShapeNet datasets, in which the method decomposes scenes into reasonable parts, outperforming existing state-of-the-art approaches.
-Index Terms:
-Part-aware reconstruction, Hybrid representation, 2D Gaussians, Superquadrics.
-{strip}[Uncaptioned image]
+Our method simultaneously optimizes superquadric meshes and Gaussians by coupling their parameters within our hybrid representation.
+On one hand, this hybrid representation inherits the advantage of superquadrics to represent different shape primitives, supporting flexible part decomposition of scenes.
+On the other hand, 2D Gaussians capture complex texture and geometry details, ensuring high-quality appearance and geometry reconstruction.
+Our method is fully unsupervised and outperforms existing state-of-the-art approaches in extensive experiments on DTU, ShapeNet, and real-life datasets.
+1 Introduction
 Figure 1: We propose a pipeline to reconstruct the scene in a part-aware manner.
@@
 Here, EMS [22] takes point cloud inputs and reconstructs non-textured primitives.
-I Introduction
 3D reconstruction from multi-view images is a long-studying problem in 3D vision and computer graphics.
-Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes, which is different from human perceptions.
-Humans understand 3D scenes or objects as different semantic parts [3].
-For example, when we see a scene, we are building high-level structural information such as scene graphs in our minds, not point clouds or detailed geometry.
-Thus, we propose to learn part-aware scene reconstruction to decompose the scenes into different individual semantic parts or shapes, which would be more helpful for tasks including scene manipulation/editing, scene graph generation, etc.
+Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes, which differ from human perception.
+Humans understand 3D scenes as different semantic parts [26].
+For instance, when observing a scene, we mentally construct high-level structural information, such as scene graphs, rather than point clouds or detailed geometry.
+Thus, we propose to learn part-aware scene reconstruction to decompose scenes into different individual semantic shapes or parts, facilitating tasks including physical simulation, manipulation/editing, 3D content generation, etc.
 There are prior methods [38, 51, 2, 28, 20, 44, 22, 23, 36, 42, 48] working on this topic, while they mainly learn the goal with 3D supervisions and cannot preserve accurate geometries, which leaves inconvenience to apply in realistic scenarios.
-While neural radiance field (NeRF) [14] has demonstrated the potential of reconstructing textured 3D scenes from multi-view images, some approaches [15, 16] try to learn part-aware objects via NeRFs.
-For example, Part-NeRF [15] represents an object as multiple neural radiance fields.
-However, the complex composition of multiple neural fields increases learning difficulty, resulting in subpar rendering outcomes and structurally insignificant information.
-Another state-of-the-art approach DBW [2] decomposes the scenes into multiple blocks represented by superquadric primitives [6], optimizing their parameters and UV texture maps to minimize rendering loss.
-Although they succeed in decomposing 3D scenes into reasonable parts, learning accurate detailed geometry and texture maps still proves challenging, or not included in the pipelines.
-Segmenting the scene into reasonable parts while preserving high-quality geometry details and textures is essential for good geometry reconstruction and rendering quality.
-To tackle the problem, we propose integrating superquadrics and 2D Gaussians [17] to reconstruct part-aware and well-textured scenes, culminating in a hybrid representation.
-The hybrid representation retains the primitive property of superquadrics, where we assume each part should be a basic shape, and shapes that cannot be fitted as a superquadric should be segmented into multiple parts.
-On the other hand, Gaussian splatting, renowned for its superior rendering capabilities in recent studies [18, 19, 20] and efficient training, is incorporated to capture intricate texture details, speeding up the reconstruction and improving the rendering quality significantly.
-Specifically, in the hybrid representation, Gaussian ellipses are attached to the surface of each superquadric mesh.
-We adopt 2D Gaussians and share their pose parameters with the attached superquadric mesh.
+While NeRF [25] has demonstrated great potential in reconstructing textured 3D scenes from multi-view images, some approaches [37, 49] try to learn part-aware objects via NeRFs.
+In particular, Part-NeRF [37] represents an object as multiple neural radiance fields.
+However, the complex composition of multiple implicit fields increases learning difficulty, resulting in subpar rendering quality and insufficient semantic decomposition.
+Another state-of-the-art approach DBW [27] decomposes scenes into blocks represented by superquadric primitives [2], optimizing their parameters and UV texture maps by minimizing rendering loss.
+Although DBW effectively decomposes 3D scenes into reasonable parts, learning accurate part-aware geometry and appearance still proves challenging as shown in Fig.
+1.
+To tackle the problem, we propose PartGS, a part-aware hybrid representation that integrates 2D Gaussians [16] and superquadrics [2] to reconstruct scenes that are both well-textured and geometrically accurate.
+There are previous approaches [15, 13, 39] combining mesh reconstruction and Gaussian splatting, where they first reconstruct the mesh and then attach Gaussians to it.
+Their purpose is to get both representations, while our goal is to get part segmentation and reconstruction at the same time.
+Our method directly learns superquadric mesh and Gaussians simultaneously by coupling parameters between them while constraining each part to be represented as a single superquadric.
+The rendering capability of coupled Gaussians drives the optimization of hybrid representation, enabling the use of their convexity to assist in the decomposition of shapes.
+The unsupervised training of part-aware reconstruction is built upon the assumption that each functional part of most objects should be a basic shape and can be represented by a superquadric.
+On the other hand, Gaussian splatting, renowned for its superior rendering capabilities and efficient training[18, 11, 24], is incorporated to capture intricate texture details, speeding up reconstruction and improving rendering quality significantly.
+In the hybrid representation, 2D Gaussians are attached to the surface of superquadrics, with their pose and shape parameters coupled to the corresponding superquadric mesh.
 Compared to the vanilla Gaussian [18] that fills up the whole occupancy with individual-parameter 3D Gaussians, our method is more compact and efficient.
-In this way, the parameters to optimize include global parameters controlling the shape, pose, and opacity of the hybrid block, as well as local parameters of spherical harmonic coefficients of the Gaussian ellipses.
-In the training stage, the parameters are optimized mainly by re-rendering loss without additional supervision.
-We observe that image rendering loss tends to settle into local minima during the deformation of superquadrics.
+The optimized parameters include global ones controlling the overall shape, pose, and opacity of hybrid blocks, and local parameters for spherical harmonic coefficients of the Gaussians.
+During training, the parameters are optimized by rendering loss without additional supervision.
+We observe that image rendering loss tends to settle into local minima during the optimization of superquadrics.
 To address this, we design several regularizers to enforce global consistency between the 3D representation and the input 2D information.
 This allows us to segment parts in a fully unsupervised manner, achieving block-level reconstruction.
-At last, to model the accurate geometry, we conduct a point-level refinement, freeing the 2D Gaussians to get out of the surface, thereby improving the modeling of irregular shapes.
-We summarize the contributions as follows:
-  • A novel hybrid representation is introduced to learn part-aware 3D scenes, where primitives model the parts and Gaussians model the textures.
+At last, to model the accurate geometry, we conduct a point-level refinement, freeing 2D Gaussians to get out of the surface, thereby improving the modeling of irregular shapes.
+Contributions are summarized as follows:
+  • A novel hybrid representation is introduced to learn part-aware textured 3D scenes.
 The hybrid representation combines the benefits of both superquadrics and Gaussian splitting, enabling accurate part reconstruction and efficient high-quality rendering.
-  • An end-to-end unsupervised pipeline is proposed for block-level and point-level part-aware reconstruction, with novel regularizers to optimize the superquadric meshes and 2D Gaussians simultaneously.
-  • Extensive experiments on public datasets demonstrate the superiority of the proposed method in part-aware reconstruction, working for part segmentation of objects and scenes, outperforming several state-of-the-art methods.
-Figure 2: Overview of the proposed method.
-It takes multi-view images as input to learn a parametric hybrid representation of superquadrics and 2D Gaussians.
+  • An end-to-end unsupervised pipeline is proposed for block-level and point-level part-aware reconstruction, with superquadric meshes and 2D Gaussians optimized simultaneously by novel regularizers.
+  • Extensive experiments on public datasets demonstrate the superiority of the proposed method in part-aware reconstruction, enabling applications of part augmentation and physical interactions.
+Figure 2: Overview of PartGS.
+PartGS takes multi-view images as input to learn a parametric hybrid representation of superquadrics and 2D Gaussians.
 The representation initializes from random superquadrics and is gradually optimized during training to get a block-level reconstruction.
-At last, we free the constraints of Gaussians to model the detailed geometry to get a point-level reconstruction.
-II Related Works
-In this section, we first introduce classic multi-view reconstruction techniques.
-Next, we provide an overview of the advancements in primitive-based 3D representation.
-Additionally, we discuss related works on surface-aligned Gaussian splatting, which inspired our hybrid representation.
-II-A Multi-View Reconstruction
-Before the era of deep learning, multi-view stereo-based methods [21, 22, 23, 24] adopt a standard pipeline that consisting of image feature matching, depth prediction, and fusion from different viewpoints.
-Recent neural approaches [25, 26, 27] combining implicit representation and neural rendering outperforms previous methods in high-fidelity reconstruction.
-These techniques typically represent surfaces implicitly via a multi-layer perceptron (MLP) and extract surfaces post-training using the Marching Cubes algorithm [28].
-Additionally, some works [29, 30] use volumetric representations to achieve fast learning.
-3D Gaussian Splatting (3D-GS) [18] was introduced as a new technique integrating point-based rendering and splitting.
-Due to its rendering speeds and quality never seen before, Gaussian splatting has sparked a research boom and has been applied in 3D reconstruction [17, 31, 32].
-Our work can be seen as an end-to-end approach to part-aware multi-view reconstruction, whose goal is to output the 3D reconstruction of parts from calibrated multi-view images.
-II-B Primitive-based 3D Representation
+Then, we free the constraints of Gaussians to model detailed geometry to get a point-level reconstruction.
+The last column demonstrates one application of PartGS: physics simulations, where one part is picked up from the scene and thrown forward.
+Please refer to the supplementary materials for more video demos.
+2 Related Works
+2.1 Primitive-based 3D Representation
 Learning structured shape representations aims to decompose objects or scenes into semantically coherent geometric primitives, providing a crucial pathway to shape understanding and generation [43, 50, 14, 9, 41, 35].
 The earliest classical approaches, dating back to the last century, such as Blocks World [32] and Generalized Cylinders [3] emphasize compactness as the central goal.
-However, these ideas struggle to handle the complexity of real-world data.
 A substantial amount of prior work has focused on processing 3D data inputs, such as point clouds, meshes, and voxel grids.
 These methods decompose the 3D data into various types of primitive ensembles, including cuboids [38, 31], superquadrics [29, 30, 44, 22], and convex shapes [8, 10].
-Recent advancements include MonteBoxFinder [41], which utilizes a RANSAC-based [46] approach with cuboids to represent the input point cloud, combined with Monte Carlo Tree Search for optimal selection.
-Alternatively, EMS (Expectation, Maximization, and Switching) [1] employs superquadrics to represent geometric primitives and proposes a probability-based approach to recover these primitives from the point cloud.
-However, such methods often rely on perfect 3D data, rarely available in real-world scenarios.
-In practice, 3D data is frequently incomplete or noisy, especially when reconstructed from 2D images.
-This noise and missing data significantly hinder the practical application of these methods.
-There are only a few studies [15, 2, 16, 47] that attempted to create structure-aware 3D representations directly from images without relying on 3D supervision.
-For example, PartNeRF [15] uses ellipses to achieve shape abstraction, modeling each part separately using a neural radiation field.
-However, using ellipses makes PartNeRF more likely to decompose shapes based solely on spatial locations, which limits its ability to recover meaningful structural information.
-ISCO [47] and DBW [2] propose decomposing shapes using 3D superquadrics, which allows for meaningful shape decomposition.
-Nonetheless, the simplicity of their shape parameters struggles to capture complex geometries, leading to poor reconstruction accuracy and challenges in rendering high-quality novel view synthesis.
-Recently, DPA-Net [16] has made great progress on 3D shape abstraction from sparse views.
-However, this approach generates redundant parts and struggles with realistic texture rendering.
-Our approach introduces a hybrid representation to learn part-aware 3D scenes while maintaining accurate structural reconstruction and efficient, high-quality rendering capabilities.
-II-C Surface-Aligned Gaussian Splatting
-Gaussian splatting [18] has been rapidly adopted in multiple fields because of its remarkable capability in rendering.
-However, the challenge of unconstrained Gaussians resulting in inaccurate geometry during reconstruction has prompted several studies [48, 49, 50, 51, 17] to align Gaussians with scene surfaces.
-Among them, SuGaR [48] addresses this by using flat 3D Gaussians to enforce the alignment with the scene surface during optimization, minimizing the difference between the signed distance functions (SDF) of the desired Gaussians and the actual Gaussians.
-GaMeS [51] introduces a hybrid representation of Gaussians and mesh, where Gaussians are attached to the triangular facets of the mesh.
-It allows for mesh-like adaptation and modification of Gaussian components.
+MonteBoxFinder [31], uses a RANSAC-based [12] with cuboids to represent point clouds, combined with Monte Carlo Tree Search for optimal selection.
+Alternatively, EMS [22] employs superquadrics to represent geometric primitives and proposes a probability-based approach to recover primitives.
+These methods rely on perfect 3D data, which is rarely available due to incompleteness or noise.
+Few studies [37, 27, 49, 1] have attempted to create structure-aware 3D representations directly from images.
+PartNeRF [37] uses ellipses with neural radiance fields, but this spatial-based decomposition limits its ability to capture meaningful structural information.
+ISCO [1] and DBW [27] use 3D superquadrics for shape decomposition, which enables more meaningful structural separation.
+However, their simple shape parameters struggle to capture complex geometries, leading to poor geometry and appearance reconstruction.
+DPA-Net [49] has advanced 3D shape abstraction from sparse views but generates redundant parts and struggles with realistic texture rendering.
+Our approach introduces a hybrid representation that learns part-aware 3D scenes while ensuring accurate structural reconstruction and efficient, high-quality rendering.
+2.2 Mesh-based Gaussian Splatting
+Gaussian splatting(GS) [18] has been rapidly adopted in multiple fields due to its remarkable rendering capability.
+Several studies [15, 7, 13, 39] attempt to align Gaussians with mesh surfaces for easier editing and animation.
+Among them, SuGaR [15] uses flat 3D Gaussians to enforce the alignment with the scene surface during optimization, minimizing the difference between the signed distance functions (SDF) of the desired Gaussians and actual Gaussians.
+GaMeS [39] introduces a hybrid representation of Gaussians and mesh, where Gaussians are attached to triangular facets of the mesh.
 Similarly, Gao et al.
-[50] propose a mesh-based Gaussian Splatting to achieve large-scale deformation effects on objects.
-NeuSG [49] combines 3DGS with implicit SDF networks, optimizing the Gaussian components and the network parameters jointly.
-Recently, 2DGS [17] proposes using 2D Gaussians instead of 3D Gaussians, leveraging 2D surface modeling to significantly enhance the geometric quality of scenes.
-Inspired by them, our work proposes a representation that combines 2D Gaussians with superquadrics, ensuring that the Gaussians conform to the surface of the superquadrics.
-III Method
-Given a set of calibrated multi-view images, we aim to learn the part-aware 3D geometry and texture information based on Gaussian splatting.
-We observe that a part-aware scene can intuitively be regarded as a composition of multiple basic shapes (geometric primitives).
-Therefore, we design a primitive-based hybrid representation to describe the 3D scene, as introduced in Section III-A.
-In Section III-B, this hybrid representation can be optimized to fit primitives as different parts, as stage 1.
-In stage 2, we refine the positions of 2D Gaussians to model the precise geometry of the original shape.
-III-A Hybrid Representation
+[13] proposes a mesh-based GS to achieve large-scale deformation effects on objects.
+Recently, 2DGS [16] proposes 2D Gaussians for surface modeling and significantly enhances the geometric quality.
+Inspired by them, we propose a representation that combines 2D Gaussians with superquadrics.
+A key distinction between our method and previous mesh-based GS approaches is that these methods require first reconstructing the scene’s mesh and then binding Gaussians to the mesh surface, which results in a non-continuous mesh.
+In contrast, our method directly optimizes the mesh through a rendering loss, enabling part-aware mesh reconstruction.
+3 Method
+Given a set of calibrated multi-view images, our goal is to learn part-aware 3D geometry and appearance.
+A part-aware reconstruction is often represented as a composition of multiple basic shapes (geometric primitives).
+We propose a primitive-based hybrid representation in Sec.
+3.1.
+In Sec.
+3.2, we demonstrate how this hybrid representation is optimized to fit primitives as parts (block-level).
+In order to further detail the geometry of each part, the parameters of Gaussians are further refined to get a point-level reconstruction.
+3.1 Hybrid Representation
 As shown in Fig.
-2, the proposed hybrid representation combines superquadrics with 2D Gaussian ellipses in a compact form.
-Superquadrics are efficient for modeling 3D scenes due to their primitive properties, but they fall short in capturing detailed geometry and texture information.
-Conversely, the 3D/2D Gaussian representation [18, 17] excels in photorealistic rendering but is cumbersome and lacks structural awareness.
-To leverage the strengths of both representations, we attach 2D Gaussian ellipses to the surface of the superquadric mesh.
-This approach retains the superquadric’s advantage of effectively parsing a 3D scene into distinct parts.
-At the same time, the Spherical Harmonics of 2D Gaussians enable the rendering of complex textures by Gaussian splatting, addressing the limitations in texture learning.
-By sharing the pose parameters between superquadrics and 2D Gaussians, we achieve a more efficient representation.
-III-A1 The Parametric Model
-The proposed hybrid model is parametric, controlled by primitive parameters and Gaussian parameters.
+2, the proposed hybrid representation combines superquadrics with 2D Gaussians in a compact form.
+To leverage the strengths of both, we attach Gaussians to the surface of superquadric meshes.
+This representation retains the superquadric’s ability to effectively parse a 3D scene into distinct parts.
+Meanwhile, spherical harmonics of Gaussians enable complex texture rendering via Gaussian splatting, addressing the texture learning limitations of previous work [22, 31, 27, 48].
+Sharing pose parameters between superquadrics and 2D Gaussians further enhances the representation’s efficiency.
+The proposed hybrid model is parametric, controlled by both primitive and Gaussian parameters, which are optimized simultaneously.
 Considering a 3D scene $\mathcal{S}$, we parse it into multiple superquadrics, i.e.
-blocks, in the first stage.
+blocks.
 Each block is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$.
 Here, $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks.
-The hybrid blocks are initialized using manually designed parameters that control pose, opacity, scale, shape, and texture.
-These parameters are optimized through a multi-view rendering loss to learn the scene.
+The hybrid blocks are defined using manually designed parameters that control pose, opacity, scale, shape, and texture.
+These parameters are optimized via differentiable rendering to parse the target scene.
 Shape and Scale Parameters.
-For each hybrid block $\mathcal{B}_{i}$, its shape is defined by the superquadric mesh.
-As introduced in [6], a superquadric comprises a series of parametric surfaces controlled by several continuous parameters.
-More concretely, there are 2 parameters $\epsilon_{1},\epsilon_{2}$ defining the superquadric mesh and 3 anisotropy parameters $s_{1},s_{2},s_{3}$ scaling the 3D axes.
-Similar to the position of each vertex of an icosphere, we can acquire the 3D position of the superquadric vertex by:
-$\scriptstyle\mathbf{v}=[s_{1}\cos^{\epsilon_{1}}(\theta)\cos^{\epsilon_{2}}(\varphi);s_{2}\sin^{\epsilon_{1}}(\theta);s_{3}\cos^{\epsilon_{1}}(\theta)\sin^{\epsilon_{2}}(\varphi)],$ (1)
-where $\theta$ and $\varphi$ are the azimuth and elevation defined in the spherical coordinate system.
-The shape and scale parameters are optimized during the training stage, thus controlling the block deformation to learn part-aware information.
+For each hybrid block $\mathcal{B}_{i}$, its geometry is controlled by superquadric parameters [2].
+More concretely, there are two shape parameters $\epsilon_{1},\epsilon_{2}$ that define the superquadric mesh, along with three scale parameters $s_{1},s_{2},s_{3}$, which scale the 3D axes.
+Similar to the positioning of vertexes in an icosphere, the 3D position of a superquadric vertex can be obtained by:
+$\footnotesize\mathbf{v}=[s_{1}\cos^{\epsilon_{1}}(\theta)\cos^{\epsilon_{2}}(\varphi);s_{2}\sin^{\epsilon_{1}}(\theta);s_{3}\cos^{\epsilon_{1}}(\theta)\sin^{\epsilon_{2}}(\varphi)],$ (1)
+where $\theta$ and $\varphi$ represent the azimuth and elevation defined in the spherical coordinate, respectively.
+The shape and scale parameters are optimized during training, thus controlling the block deformation to learn part-aware geometry.
 Pose Parameters.
-The initial hybrid blocks are randomly sampled and disorganized in the space, so it is necessary to align their poses to fit within the 3D scene.
-The vertex position acquired in Eq.
-1 is in the local coordinate system of the hybrid block, so we need to use the pose parameters to transform the position into the world space.
-We define the pose parameter of the $i$-th hybrid block as the combination of rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$.
-Hence, the vertex position in the world space is calculated by: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $i,j$ denotes the indexes of block and vertex respectively.
-In general, the Gaussians placed on the superquadric surface also have their individual poses, which causes a large amount of parameters.
-To address this issue, we propose to share the pose of the superquadric with the Gaussians.
-Given a posed superquadric, we directly determine the center points of the Gaussians through uniform sampling on the triangular faces and let the normal of 2D Gaussians be the same as the faces.
-Inspired by GaMeS [51], we assign the rotation matrix $\mathrm{R}_{v}$ and scaling $\mathrm{S}_{v}$ of Gaussians based on the vertices’ positions.
-For a given triangular face $V$ and its three vertices $v_{1}$, $v_{2}$, $v_{3}\in\mathbb{R}^{3}$, we define the orthonormal vectors $r_{1}$, $r_{2}$, $r_{3}\in\mathbb{R}^{3}$ to construct the rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$.
-The first vector $r_{1}$ is aligned with the normal vector of $V$ to ensure the Gaussians align with the triangular surface.
+The pose of the $i$-th hybrid block is defined by its rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$.
+The vertex position from Eq.
+1 is transformed from the local coordinate to world space as: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $j$ indexes the vertices.
+In prior methods such as SuGaR [15] and GaMeS [39], Gaussians are placed on the reconstructed mesh, and they have individual poses and shapes for better appearance modeling.
+Differently, in our method, the role of Gaussians is to construct a differentiable rendering that bridges images and superquadrics.
+To achieve this, we couple the parameters of Gaussians with superquadrics.
+Given a posed superquadric, centers of Gaussians are determined by uniformly sampling on triangular faces, and poses of Gaussians are determined by face vertices.
+Inspired by GaMeS [39], we assign the rotation matrix $\mathrm{R}_{v}$ and scaling $\mathrm{S}_{v}$ of Gaussians based on vertices’ positions.
+For a given triangular face $V$ and its vertices $v_{1}$, $v_{2}$, $v_{3}\in\mathbb{R}^{3}$, we define orthonormal vectors $r_{1}$, $r_{2}$, $r_{3}\in\mathbb{R}^{3}$ to construct the rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$.
+The vector $r_{1}$ is aligned with the normal vector of $V$ to ensure Gaussians align with the triangular surface.
 The vector $r_{2}$ is defined as the vector from the center $m=\text{mean}(v_{1},v_{2},v_{3})$ of the triangle to the first vertex $v_{1}$:
@@
 $r_{3}=\frac{{ort}(v_{2}-m;r_{1},r_{2})}{||{ort}(v_{2}-m;r_{1},r_{2}||},$ (3)
-where ${ort}$ represents the orthonormalization step in the Gram-Schmidt [52]. For the scaling parameters of the 2D Gaussians, we use:
+where ${ort}$ represents the orthonormalization in Gram-Schmidt [4]. For the scaling of 2D Gaussians, we use:
 $\mathrm{S}_{v}=\text{diag}(s_{v}^{2},s_{v}^{3}),$ (4)
-where $s_{v}^{2}=c||m-v_{1}||$, and $s_{v}^{3}=c<v_{2},r_{3}>$, and $c$ is a size control hyperparameter. We place a fixed number of $n$ Gaussians on each triangular face.
-This design avoids the additional cost of optimizing the pose and covariance matrix of Gaussians, making it more efficient for representing the scene.
-The pose parameters are optimized via gradient descent so that the hybrid blocks can fit the scene.
+where $s_{v}^{2}=c||m-v_{1}||$, $s_{v}^{3}=c<v_{2},r_{3}>$, and $c$ is a size control hyperparameter.
+We place a fixed number of Gaussians on each triangular face.
+These designs eliminate the need to learn the geometry of Gaussians, focusing solely on their texture, thereby enhancing the efficiency of the representation.
 Opacity Parameters.
 We define the total number of hybrid blocks as $M$.
-However, in practice, a typical scene does not contain exactly $M$ blocks, so we retain only the meaningful ones.
-Therefore, we utilize a learnable parameter $\tau_{i}$ to represent each block’s opacity.
-In the optimization process, only blocks with $\tau_{i}$ greater than a certain threshold are retained.
+However, in practice, a typical scene does not contain exactly $M$ blocks.
+Therefore, only the meaningful blocks are retained.
+To achieve this, we introduce a learnable parameter $\tau_{i}$ to represent each block’s opacity.
+During the optimization process, only blocks with $\tau_{i}$ greater than a certain threshold are retained.
 Note that Gaussians on the same block share the same $\tau$ for opacity in rasterization.
 Texture Parameters.
-The texture information is modeled using 2D Gaussians positioned on the surface of the superquadric mesh.
-Spherical harmonics control the texture of each Gaussian ellipse.
-Therefore, we define the coefficients of the spherical harmonics as texture parameters and optimize them for rendering view-dependent images.
-III-A2 Rendering
-With Gaussians attached to the surface of each block, we achieve view-dependent rendering through tile-based rasterization as Gaussian Splatting [18, 17].
-Given a view, Gaussians on the superquadric surface are projected onto the image space, forming an RGB image.
-Initially, the screen space determines the bounding box for each Gaussian.
-Subsequently, these Gaussian ellipses are sorted according to their depths of center to the image plane.
-Finally, volumetric alpha compositing [53] is utilized to integrate the alpha-weighted RGB values for each pixel.
-To formulate the process, considering a viewing transformation ${W}$, the covariance matrix ${\Sigma}_{i}^{{}^{\prime}}$ of $i$-th Gaussian in the camera coordinate system is calculated by:
-${\Sigma}_{i}^{{}^{\prime}}={J}{W}{\Sigma}{W}^{T}{J}^{T},$ (5)
-where ${J}$ is the Jacobian of the affine approximation of the projective transformation, and ${\Sigma}$ is the covariance matrix of the Gaussian ellipse.
-Note that the last row and column of ${\Sigma}$ are omitted since we adopt 2D Gaussians.
-Following alpha compositing, we first calculate an alpha value for each Gaussian ellipse:
-$\alpha_{i}=\tau_{i}\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{u}_{i})^{T}{\Sigma_{i}^{{}^{\prime}}}_{-1}(\mathbf{x}-\mathbf{u}_{i}))).$ (6)
-Here, $\mathbf{u}_{i}$ is the center coordinate of the projected Gaussian ellipse, and $\tau_{i}$ is the opacity of the block where the $i$-th Gaussian is located.
-The calculated alphas are sorted according to their depths from the image plane.
-Meanwhile, we can acquire the color value $c_{i}$ from the spherical harmonics of $\mathcal{N}$ ordered points, thus obtaining the rendering RGB value:
-$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (7)
-This process is differentiable and can optimize the hybrid representation through gradient descent, and the optimization/learning is introduced in Sec. III-B.
-III-B Optimizing the Hybrid Representation
-In this section, we introduce how to optimize the hybrid representation.
-By minimizing the rendering loss across multi-view images only, we found it unstable to position the hybrid blocks (primitives) accurately.
-Thus, we introduce several regularizations and perform two optimization stages (block-level and point-level optimizations).
-III-B1 Block-level Optimization
-Besides the image rendering loss, we employ regularizers consisting of coverage, overlap, parsimony, and opacity entropy.
-The coverage loss encourages hybrid blocks to cover only the meaningful regions; the overlap loss prevents different blocks from overlapping; the parsimony loss regularizes the number of existing blocks; the opacity entropy encourages opacities of hybrid blocks close to binary.
-Rendering Loss.
-Based on Eq.
-7, we can obtain the rendered image $\hat{\mathbf{I}}$ associated with a given camera pose.
-We calculate the rendering loss between $\hat{\mathbf{I}}$ and the ground truth image $\mathbf{I}$ by associating $L_{1}$ with a D-SSIM term[18]:
-$L_{\text{ren}}=(1-\lambda)\ L_{1}+\lambda L_{\text{D-SSIM}}.$ (8)
+The texture is modeled using 2D Gaussians positioned on the surface of the superquadrics.
+Spherical harmonics of each Gaussian control the texture and are optimized for rendering view-dependent images [18].
+3.2 Optimizing the Hybrid Representation
+This section describes the optimization of the hybrid representation.
+We found that minimizing the rendering loss across multi-view images alone led to instability in positioning hybrid blocks (primitives) accurately.
+Therefore, we introduce several regularizations and perform two optimization stages: block-level and point-level optimization.
+3.2.1 Block-level Optimization
+In addition to the image rendering loss, we employ regularizers consisting of coverage, overlap, parsimony, and opacity entropy.
+The coverage loss encourages hybrid blocks to cover only meaningful regions; the overlap loss prevents blocks from overlapping; the parsimony loss regularizes the number of existing blocks; and the opacity entropy encourages block opacities close to binary.
+Rendering Loss. Based on the rendering process of Gaussian spaltting [18], the rendering loss is calculated by associating $L_{1}$ with a D-SSIM term[18] :
+${\mathcal{L}}_{\text{ren}}=(1-\lambda)\ L_{1}+\lambda L_{\text{D-SSIM}}.$ (5)
 Coverage Loss.
-This loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending outside the object.
-To determine the occupancy relationship in 3D space, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=1-\Psi_{i}(x)$.
-Here, $\Psi_{i}$ denotes the superquadric inside-outside function [6] associated to the block $i$.
-Consequently, $D_{i}(x)\geq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)<0$ if $x$ is outside of the block.
-Similar to NeRF [14], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses.
-Given the object mask, we associate each ray $r\in\mathcal{R}$ with a binary label, such that $l_{r}=1$ if the ray $r$ is inside the mask, otherwise $l_{r}=0$.
+This loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending beyond the object’s boundaries.
+To determine the 3D occupancy field based on the blocks, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=1-\Psi_{i}(x)$.
+Here, $\Psi_{i}$ denotes the superquadric inside-outside function [2] associated with the block $i$.
+Consequently, $D_{i}(x)\geq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)<0$ if $x$ is outside the block.
+Inspired by NeRF [25], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses.
+Given the object mask, we associate each ray $r\in\mathcal{R}$ with a binary label: $l_{r}=1$ if the ray $r$ is inside mask, otherwise $l_{r}=0$.
 The coverage loss is defined as:
 ${\mathcal{L}}_{\text{cov}}({\mathcal{R}})=\sum_{r\in{\mathcal{R}}}l_{r}L_{\text{cross}}(r)+(1-l_{r})L_{\text{non-cross}}(r).$ (6)
-Here, $L_{\text{cross}}$ encourages rays to intersect with blocks , while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks.
-Intuitively, this means that there is at least one in sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not lie within any block.
-This process can be expressed as:
+Here, $L_{\text{cross}}$ encourages rays to intersect with blocks, while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks. They are defined as:
 $L_{\text{cross}}(r)=\text{ReLU}\Big{(}\min_{i\in\mathcal{M}}\min_{x_{j}^{r}\in{\mathcal{X}}_{r}}D_{j}(x_{j}^{r})\Big{)},$ (7)
@@
 where $x_{j}^{r}$ means the $j$-th sampled points along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$.
+Intuitively, this means that there is at least one in sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not lie within any block.
 Overlap Loss.
-We add a regularization term to encourage individual blocks to be non-overlapping.
-Considering the difficulty in directly calculating the overlap of locks, we use a Monte Carlo method similar to  [2].
-Practically, we sample multiple 3D points in space and penalize each point for being inside more than $k$ blocks.
+We add a regularization term to encourage individual blocks to remain non-overlapping.
+Given the difficulty of directly calculating block overlap, we use a Monte Carlo method similar to  [27].
+Practically, we sample multiple 3D points in space and penalize each point that lies inside more than $k$ blocks.
 Based on the definition of the superquadric inside-outside function, we devise a soft occupancy function:
 $\mathcal{O}_{i}^{x}=\tau_{i}\text{ sigmoid}(\frac{1-\Psi_{i}(x)}{\gamma}),$ (9)
-where $\gamma$ is a temperature parameter. Therefore, given a set of $N$ sampled 3D points $\Omega$, the overlap loss can be expressed as:
-$L_{\text{over}}=\frac{1}{N}\sum_{x\in\Omega}\text{ReLU}\left(\sum_{i\in\mathcal{M}}\mathcal{O}_{i}^{x}-k\right).$ (13)
-Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we add a loss term that regularizes the number of active blocks. This loss is defined as:
-$L_{\text{par}}=\frac{1}{M}\sum_{i\in\mathcal{M}}\sqrt{\tau_{i}}.$ (14)
+where $\gamma$ is a temperature parameter. Therefore, given a set of $N$ sampled 3D points $\Omega$, the overlap loss is expressed as:
+${\mathcal{L}}_{\text{over}}=\frac{1}{N}\sum_{x\in\Omega}\text{ReLU}\left(\sum_{i\in\mathcal{M}}\mathcal{O}_{i}^{x}-k\right).$ (10)
+Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we add a loss term that regularizes the block opacity ($\tau$). This loss is defined as:
+${\mathcal{L}}_{\text{par}}=\frac{1}{M}\sum_{i\in\mathcal{M}}\sqrt{\tau_{i}}.$ (11)
 Opacity Entropy Loss.
-Naturally, during optimization, the opacity of blocks inside the object region tends to approach 1, while the opacity of blocks outside the object region tends to be 0.
+Naturally, during optimization, the opacity of the block inside the object region tends to approach 1, while the opacity of the block outside the object region tends to approach 0.
 To facilitate this, we associate the opacities of the blocks with the labeled rays.
-Specifically, we employ a cross-entropy loss, $L_{ce}$, between the opacity and mask labels as follows:
-${\mathcal{L}}_{\text{opa}}({\mathcal{R}})$ $=\frac{1}{|{\mathcal{R}}|}\sum_{r\in{\mathcal{R}}}L_{ce}\left(\max_{i\in\mathcal{M}}\tau_{i}(x^{r}),l_{r}\right)~{},$ (15)
-$\left\{x^{r}\in{\mathcal{X}}_{r}\mid D_{i}(x^{r})\leq 0\right\}.$
-Thus, the total loss is the weighted summation of the above terms:
-$L=L_{\text{ren}}+\lambda_{\text{cov}}L_{\text{cov}}+\lambda_{\text{over}}L_{\text{over}}+\lambda_{\text{par}}L_{\text{par}}+\lambda_{\text{opa}}L_{\text{opa}}.$ (16)
+Specifically, we use a cross-entropy loss $L_{ce}$, between the opacity and mask labels as follows:
+${\mathcal{L}}_{\text{opa}}({\mathcal{R}})=\frac{1}{|{\mathcal{R}}|}\sum_{r\in{\mathcal{R}}}L_{ce}\left(\max_{i\in\mathcal{M}}\tau_{i}(x^{r}),l_{r}\right)~{}.$ (12)
+Here, only points $x^{r}$ inside the blocks are sampled, e.g., $\left\{x^{r}\in{\mathcal{X}}_{r}\mid D_{i}(x^{r})\leq 0\right\}$.
+The total loss is the weighted summation of the loss terms described above:
+${\mathcal{L}}={\mathcal{L}}_{\text{ren}}+\lambda_{\text{cov}}{\mathcal{L}}_{\text{cov}}+\lambda_{\text{over}}{\mathcal{L}}_{\text{over}}+\lambda_{\text{par}}{\mathcal{L}}_{\text{par}}+\lambda_{\text{opa}}{\mathcal{L}}_{\text{opa}}.$ (13)
 Adaptive Number of Blocks.
-Given that the number of parts varies by scene, we allow the number of blocks to adjust during optimization adaptively.
+Given that the number of parts varies across scenes, we allow the number of blocks to adjust during optimization adaptively.
 When the opacity of an existing block falls below a threshold $t$, this block is immediately removed.
 Additionally, we explore a block-adding mechanism using point clouds that are randomly generated within the scene area.
-Specifically, we employ DBSCAN [54] to cluster point clouds within the perspective of the object masks and not covered by any blocks.
-A new primitive is then introduced at the center of each point cloud cluster.
+Specifically, we use DBSCAN [34] to cluster initial point clouds (e.g., accessible by-products of COLMAP [33]) that are not covered by any blocks.
+A new block is then introduced at the center of each point cloud cluster.
 The other parameters of these new blocks are initialized randomly.
 At last, block-level reconstruction gets a 3D scene composited by several superquadrics, as shown in Fig. 2.
-III-B2 Point-level Optimization
-The primitive-based hybrid representation effectively decomposes the shape into parts but underperforms with complex objects.
-To address this, we perform a refinement stage for precise geometry fitting.
-In this stage, we free the constraint that Gaussian ellipses must be at the face of the superquadric mesh, allowing Gaussians to move freely to fill in the complex regions.
-Moreover, to prevent Gaussian ellipses from one block passing through another, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
-$L_{\text{enter}}=\frac{1}{N}\sum_{x\in\Omega}\sum_{m\in\mathcal{M}\setminus\{\delta\}}\text{ReLU}(-D_{m}(x)),$ (17)
+3.2.2 Point-level Optimization
+The primitive-based hybrid representation effectively decomposes the shape into parts but performs suboptimally for complex objects.
+To address this, we introduce a refinement stage for more precise geometry fitting.
+In this stage, we decouple the constraint between Gaussians and superquadrics, enabling the Gaussians to be freely optimized.
+Furthermore, to prevent Gaussian components from one block passing through another and disrupting the continuity and plausibility of the part decomposition, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
+${\mathcal{L}}_{\text{enter}}=\frac{1}{N}\sum_{x\in\Omega}\sum_{m\in\mathcal{M}\setminus\{\delta\}}\text{ReLU}(-D_{m}(x)),$ (14)
 where $\Omega$ is the sampled Gaussian set, and $\delta$ represents the identifier of the block to which the Gaussian $x$ belongs.
-After this stage, as shown in Fig.
-2, the reconstruction is more precise and better fits the original shape.
-Figure 3: Qualitative comparisons on DTU [55].
-We compare our approach with state-of-the-art baselines on the DTU dataset with the background removed.
-The last two columns show our block-level and point-level reconstructions, respectively.
+As seen in Fig.
+2, the reconstruction becomes more accurate and aligns better with the target shape after this stage.
+Figure 3: Qualitative comparison on DTU [17] and ShapeNet [17] .
+The first two rows are DTU examples, and the last two are ShapeNet examples, respectively.
 Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
-Figure 4: Qualitative comparisons on ShapeNet [55].
-We compared our approach with state-of-the-art baselines across four categories.
-The last two columns display our block-level and point-level reconstructions, respectively.
-Our method uniquely provides reasonable 3D part decomposition while simultaneously capturing detailed geometry.
 Figure 4: Qualitative results on BlenderMVS [45] and self-capatured data.
-We demonstrate the RGB renderings and decomposed parts from novel views across a variety of objects.
-The first five examples are from the BlenderMVS dataset, and the remaining examples are from our own captured scenes.
+We present RGB renderings and decomposed parts from novel views.
+The top examples are from the BlenderMVS dataset, and the last example is from our captured scenes.
 Figure 5: Ablation studies on key strategies.
 The block-level visual comparisons illustrate the impact of adopting our proposed strategy.
-The first row indicates results without the strategy, while the second row shows results with the strategy implemented.
-TABLE I: Quantitative comparison on DTU [55].
-The Chamfer Distance (CD) between the 3D reconstruction and the ground-truth is reported in 15 scenes.
-The best results are bolded, and the average numbers of primitives found (#P) that are greater than 10 are underlined.
-Chamfer Distance (CD) per scene Mean Mean
-Method Input S24 S37 S40 S55 S63 S65 S69 S83 S97 S105 S106 S110 S114 S118 S122 CD #P
-EMS [1] GT point cloud 6.32 3.54 2.99 4.30 4.16 4.01 3.75 3.24 4.97 4.34 4.16 7.62 7.58 4.46 4.03 4.65 7.7
-MBF [41] GT point cloud 3.12 2.66 3.84 2.54 1.59 2.11 2.19 2.01 2.32 2.45 2.17 2.12 3.83 2.02 2.55 2.50 34.1
+The first row shows results without the strategy, and the second with the strategy implemented.
+Chamfer distance per scene Mean Mean
+Method Input Renderable S24 S37 S40 S55 S63 S65 S69 S83 S97 S105 S106 S110 S114 S118 S122 CD #P
+EMS [22] 3D GT 6.32 3.54 2.99 4.30 4.16 4.01 3.75 3.24 4.97 4.34 4.16 7.62 7.58 4.46 4.03 4.65 7.7
+MBF [31] 3D GT 3.12 2.66 3.84 2.54 1.59 2.11 2.19 2.01 2.32 2.45 2.17 2.12 3.83 2.02 2.55 2.50 34.1
 EMS [22] + Neus [40] Iamge 5.99 5.56 4.43 4.32 5.42 6.14 3.75 3.96 4.63 4.34 5.88 5.11 4.29 4.83 3.53 4.97 8.87
@@
 Ours (Point-level) Image 0.95 1.38 0.55 0.71 1.02 1.39 1.16 1.33 1.49 0.80 0.95 1.31 0.69 1.20 0.82 1.05 5.9
-TABLE II: Quantitative comparisons on DTU [55].
-We compare our method to PartNeRF, DBW, and 2DGS in terms of PSNR, SSIM, LIPPS, Chamfer Distance, and reconstruction time.
-All running time measurements were conducted on a single GTX RTX3090 GPU.
-Excluding the results of 2DGS, the best results are in bold font, and the second-best results are underlined.
-Method Structure CD $\downarrow$ PSNR $\uparrow$ SSIM  $\uparrow$ LIPPS $\downarrow$ Time $\downarrow$
-2DGS [17] 0.96 34.07 0.99 0.019 $\sim$ 18 mins
-PartNeRF [15] 9.59 17.97 0.77 0.246 $\sim$ 8 hours
-DBW [2] 4.73 16.44 0.75 0.201 $\sim$ 2 hours
-Ours (Block-level) 4.19 19.84 0.82 0.189 $\sim$ 30 mins
-Ours (Point-level) 1.05 35.04 0.99 0.015 $\sim$ 50 mins
-TABLE III: Quantitative comparison on ShapeNet [57].
-The Chamfer Distance (CD) between the 3D reconstruction and the ground-truth is reported on objects of four categories.
-The best results are bolded, and the second-best results are underlined.
+Table 1: Quantitative comparison on DTU [17].
+The Chamfer distance between the 3D reconstruction and the ground-truth is reported in 15 scenes.
+The best results are bolded, and the average numbers of primitives found (#P) that are greater than 10 are underlined.
+Method Part-aware CD $\downarrow$ PSNR $\uparrow$ SSIM  $\uparrow$ LIPPS $\downarrow$ Time $\downarrow$
+[…]
+2DGS [16] 0.81 34.07 0.99 0.019 $\sim$ 10 m
+PartNeRF [37] 9.59 17.97 0.77 0.246 $\sim$ 8 h
+DBW [27] 4.73 16.44 0.75 0.201 $\sim$ 2 h
+Ours (Block-level) 4.19 19.84 0.82 0.189 $\sim$ 30 m
+Ours (Point-level) 0.95 35.04 0.99 0.015 $\sim$ 40 m
+Table 2: Quantitative results on DTU [17].
+Our method outperforms all part-aware approaches in image synthesis quality, reconstruction accuracy, and efficiency.
+Neuralangelo’s results are from the original paper, with all times measured on an RTX 3090 GPU.
 Chamfer Distance  $\downarrow$ Primitives (#P ) Mean Mean
 Method Input Airplane Table Chair Gun Airplane Table Chair Gun CD #P
-EMS [1] GT point cloud 3.40 6.92 19.0 2.02 9.4 7.88 10.3 8.4 7.84 9.0
+EMS [22] 3D GT 3.40 6.92 19.0 2.02 9.4 7.88 10.3 8.4 7.84 9.0
+MBF [31] 3D GT 2.83 2.18 1.59 2.32 10.85 13.9 13.4 14.3 2.21 13.1
 PartNeRF [37] Image 2.29 2.77 2.30 2.46 8.0 8.0 8.0 8.0 2.46 8.0
@@
 Ours (Point-level) Image 1.29 1.72 0.94 1.07 3.9 6.6 7.6 5.0 1.25 5.8
-TABLE IV: Quantitative comparison on ShapeNet [57].
-We report the PSNR, SSIM, LIPPS, Chamfer Distance, and reconstruction time for our method, PartNeRF, as well as DBW and 2DGS across four categories.
-Excluding the results of 2dgs, the best results are bolded, and the second-best results are underlined.
+Table 3: Quantitative comparison on ShapeNet [6]. We report Chamfer distance and the number of parts. The best results are bolded, and the second-best results are underlined.
 Chamfer Distance  $\downarrow$ PSNR  $\uparrow$ SSIM  $\uparrow$ LIPPS  $\downarrow$
-Method Structure Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun
+Method Part-aware Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun
 2DGS [16] 1.47 2.37 0.50 1.03 40.89 39.80 39.05 41.74 0.994 0.990 0.990 0.994 0.009 0.027 0.017 0.009
@@
 DBW [27] 3.61 7.33 6.19 2.09 26.11 23.84 20.25 28.72 0.950 0.915 0.892 0.960 0.074 0.136 0.132 0.042
-Ours (Block-level) 2.47 2.15 2.32 1.78 26.93 26.92 23.91 27.97 0.936 0.911 0.891 0.942 0.082 0.139 0.126 0.072
+Ours (Block-level) 2.47 2.15 2.32 1.78 27.94 27.98 24.92 29.95 0.959 0.925 0.906 0.963 0.072 0.129 0.122 0.047
 Ours (Point-level) 1.29 1.72 0.94 1.07 41.18 36.80 36.07 39.51 0.992 0.973 0.977 0.989 0.014 0.070 0.038 0.021
-IV Experiments
-IV-A Experimental Settings
-Datasets.
-We conduct evaluation experiments on two public datasets: DTU [55] and ShapeNet [57].
-DTU is an MVS dataset comprising 80 forward-facing scenes, with 49-60 images captured for each scene.
-For evaluation, we utilize 15 scenes used in 2DGS [17].
-We downsample the images to a resolution of 400 $\times$ 300 for efficiency.
-To validate the effectiveness of our method in segmenting parts of man-made objects, we create a ShapeNet dataset that includes four categories: Chair, Table, Gun, and Airplane.
-Each category comprises 15 distinct objects, offering diverse instances for comprehensive evaluation.
-We randomly sample 100 camera poses on the sphere centered on each shape and render images at a resolution of 400$\times$400.
-The rendered images are equally split into training and testing sets.
-Additionally, to explore our method’s potential in handling real-life data, we present qualitative results on the BlenderMVS dataset [56] and self-captured scenes.
-For BlenderMVS data, we utilize standard camera poses.
-For self-captured scenes, we recover the camera poses using COLMAP [21] and normalize them following IDR [58].
-All images are resized to 400 $\times$ 300 pixels.
-Object masks are obtained using Grounded SAM [59].
-Baselines.
-We compare our method with four state-of-the-art works on 3D shape decomposition: EMS [1], MonteboxFinder (MBF) [41], PartNeRF [15] and DBW [2].
-MBF and EMS are applied to point clouds, utilizing cuboids and superquadrics, respectively, to fit the 3D points.
-The input point clouds are sampled from either the original 3D ground truth (GT) or the mesh reconstructed by the state-of-the-art MVS method Neus [27].
-We adhere to the testing procedure outlined by DBW [2] and randomly sample 5K and 200K points from the GT meshes as the input for EMS and MBF, respectively.
-Similar to our problem setting, PartNeRF and DBW use images directly as input to construct structural 3D representations.
-To ensure a fair comparison, we discarded the pre-defined ground plane in DBW since the test scenario consists solely of foreground objects.
-For PartNeRF, we perform instance-specific training and set the same number of parts as in our method.
-Additionally, we compare our method with 2DGS [17], the state-of-the-art Gaussian Splatting reconstruction method.
-Metrics.
-We evaluate from 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time.
-The 3D reconstruction quality is measured by the Chamfer Distance (CD) between the recovered geometry and GT, reflecting the accuracy of 3D reconstructions.
-View synthesis uses three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [60], and Learned Perceptual Image Patch Similarity (LPIPS) [61].
-Shape parsimony and reconstruction time are measured by the number of parts and running time, where fewer parts facilitate easier part editing or manipulation.
-IV-B Implementation Details
-We improve the 2DGS [17] architecture by adding a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq.
-17 during refinement optimization.
-We employ regularization in 2DGS to achieve better geometric reconstruction, including depth distortion maps, depth maps, and normal maps.
-Additionally, we introduce a mask cross-entropy to filter out extra noise Gaussians.
-To extract the meshes from 2D Gaussians, we use truncated signed distance fusion (TSDF) to fuse the rendered depth map, utilizing Open3D [62].
-In the ShapeNet benchmark, we observed a common degradation in the mesh reconstruction by 2DGS, where the meshes of objects contain many holes.
-This issue arises due to the lack of textures in ShapeNet images, causing 2DGS to use large-scale Gaussians.
-This leads to depth inconsistencies from multiview, resulting in holes in the constructed mesh.
-To address this, we apply scale regularization to limit the scale of large Gaussian components.
-We validate this design in Sec.
-IV-D.
+Table 4: Quantitative results on ShapeNet [6]. We report the Chamfer distance and novel view synthesis results across four categories.
+4 Experiments
+We evaluate our approach on 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time, comparing it with previous state-of-the-art methods.
+4.1 Datasets
+We conduct comparative experiments on two public datasets: DTU [17] and ShapeNet [6].
+To validate the effectiveness of our method in man-made objects, we construct a ShapeNet subset with four categories: Chair, Table, Gun, and Airplane.
+The DTU dataset comprises 15 commonly used scenes, and each category in the ShapeNet dataset contains 15 objects.
+Additionally, to demonstrate our method’s potential with real-world data, more qualitative results are presented on BlenderMVS dataset [45] and self-captured scenes.
+4.2 Implementation Details
+We build upon the 2DGS [16] architecture and add a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq.
+14.
+In the ShapeNet benchmark, we observe a common degradation in mesh reconstruction as shown in Fig.
+7.
+To address this, we propose Gaussian scale regularization, as validated in Sec.
+4.4.
 The same hyperparameters are used for all experiments.
 We set the initial number of primitives $M$ to 8.
-In the hybrid representation, each superquadric mesh is a level-2 icosphere (320 triangular face).
-Each triangular face contains 100 Gaussians, with a scale parameter $c$ of 0.15 and a block filter threshold $t$ of 0.2.
 The temperature parameter $\gamma$ in Eq.
@@
 We train the hybrid representation for 30k iterations, followed by refinement optimization for another 30k iterations.
-During block-level optimization, the block-adding operation is carried out at the 5000th and 10000th iterations.
-The point clouds used for block-adding are randomly initialized within the space $[-1,1]^{3}$.
-The number of sampled points in each ray is 2048 in Eq.
-9 and Eq.
-15.
+During block-level optimization, the block-adding operation is carried out at the 5k-th and 10k-th iterations.
 The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.01, and 0.01, respectively.
-IV-C Evaluations
-IV-C1 DTU Benchmark
-We utilize the official DTU Chamfer Distance (CD) evaluation [55] to compare the performance between our method and these state-of-the-art 3D decomposition techniques, as shown in Tab.
-I.
-Additionally, we present the input type and the average number of primitives used.
+For more details on the dataset, metrics, baselines, and method, please refer to the supplementary materials.
+4.3 Evaluations
+4.3.1 DTU and Shapenet Benchmark
+Geometry Reconstruction.
+In Tab.
+1 and Tab.
+3, we compare our geometry reconstruction to SOTA shape decomposition methods on Chamfer distance and training time using DTU and Shapenet dataset.
 Our method consistently outperforms in all scenes compared to prior works.
-Even without refinement optimization, our method still provides satisfactory CD errors.
-PartNeRF has the largest CD errors in the DTU dataset.
-We speculate that this is due to PartNeRF’s difficulty in handling forward scenes.
-In contrast, MBF achieves a low CD error but at the cost of using significantly more primitives, leading to over-decomposition and the inclusion of meaningless parts.
-Although EMS and DBW utilize a reasonable number of primitives, they yield higher CD errors than the proposed method.
-We compare the quality of view synthesis and reconstruction time with PartNeRF, DBW, and 2DGS in Table II.
-Even without point-level optimization, our method achieves superior image rendering quality compared to other state-of-the-art methods.
-With the addition of refinement, both image rendering quality and reconstruction accuracy are further improved, nearing the performance of the state-of-the-art Gaussian-based reconstruction method, 2DGS [17].
-However, 2DGS can only produce unstructured meshes.
-Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than the implicit reconstruction method PartNeRF and over 2X faster than DBW.
-Figure 3 presents visual comparisons of our method with existing state-of-the-art baselines.
-Even with ground-truth 3D points, EMS and MBF still struggle to extract reasonable compositions.
-PartNeRF reconstructs reasonable geometry but fails to decompose the scenes meaningful.
-DBW experiences frequent failures, particularly when parts of the geometry are not reconstructed due to the absence of primitives in these regions at initialization.
-In contrast, our approach consistently produces interpretable 3D decompositions (second-to-last column) and, through refinement, achieves more detailed structures (last column).
-IV-C2 ShapeNet Benchmark
-We evaluate our approach on four categories of ShapeNet [57].
-Since MBF [41] fails on nearly all objects of the ShapeNet benchmark, we compare only against EMS [1], PartNeRF [15] and DBW [2].
-Fig.
-4 presents the qualitative comparisons of the 3D decomposition of eight object shapes, with two instances per category.
-Tab.
-III and Tab.
-IV deliver detailed quantitative comparisons on reconstruction quality, novel view synthesis, and primitives parsimony.
-For clarity, the chamfer distance is multiplied 100 times in the Shapenet benchmark.
-EMS yields large Chamfer distances, especially in the chair and table categories, despite fitting superquadrics from GT 3D points.
-Similarly, DBW often fails to recover complete geometry in these categories, resulting in significant reconstruction errors.
-However, it performs better with guns and airplanes due to their simpler shapes.
-The non-convex nature of chairs and tables poses substantial challenges for EMS and DBW in generating reasonable geometry.
-PartNeRF benefits from NeRF’s geometric reconstruction capabilities to produce accurate geometry and good novel view renderings.
-However, it has two main disadvantages.
-On the one hand, PartNeRF cannot adaptively adjust the primitives, resulting in a fixed number of parts.
-Estimating the number of parts in advance is impractical.
-On the other hand, its use of ellipsoidal primitives limits its ability to decompose meaningful parts, such as separating the chair seat and tabletop into several components, as shown in the three columns of Fig.
+As shown in Fig.
+3, our approach consistently produces interpretable 3D decompositions with further refinement achieving more detailed geometry (the last two columns).
+MBF [31] achieves a low CD error but at the cost of using significantly more primitives, leading to over-decomposition and the inclusion of meaningless parts.
+Moreover, as shown in Tab.
+2 and Tab.
+4, our approach achieves competitive results with advanced Gaussian-based methods(e.g., 2DGS [16]) and surface reconstruction approaches (e.g., Neuralangelo [21]), which are limited to producing unstructured meshes.
+Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than PartNeRF and over 3X faster than DBW.
+Appearance Reconstruction.
+Beyond part decomposition, our method enables high-fidelity image synthesis, attributed to integrating Gaussian splatting within superquadric representations.
+EMS and MBF lack image rendering capabilities, while PartNeRF and DBW yield low-quality view synthesis.
+In contrast, our method achieves high-quality appearance rendering results, as demonstrated in Tab.
+2 and Tab.
 4.
-Conversely, the superior performance of our approach is demonstrated across all categories, visually and quantitatively.
-In addition, thanks to Gaussian’s powerful rendering capabilities, clear advantages in novel view synthesis are observed over PartNeRF and DBW, as shown in Tab.
-IV.
-The quantitative comparisons against 2DGS [17] further validate the capability of our method on reconstruction accuracy.
-Our method outperforms 2DGS in the chair and airplane categories and achieves comparable performance in the other two categories.
-TABLE V: Ablation study on the ShapeNet [57] chair benchmark.
-We report Chamfer Distance (CD), rendering metrics and the number of primitives (#P) of block-level reconstruction and point-level reconstruction.
-The best results are in bold font, and the second-best results are underlined.
+Figure 6: Qualitative comparison with SAM-based methods. Our method achieves finer-grained and more distinct part decomposition.
 Block-level Point-level
 […]
-Complete model $2.32$ $23.91$ $0.891$ $0.126$ $\underline{0.94}$ $\underline{36.07}$ $0.977$ $0.038$ $6.6$
-w/o $L_{\text{over}}$ $\underline{2.11}$ $\underline{24.13}$ $\underline{0.896}$ $\underline{0.122}$ $1.34$ $36.04$ $\underline{0.977}$ $\underline{0.037}$ $7.6$
-w/o $L_{\text{opa}}$ $5.72$ $20.12$ $0.853$ $0.174$ $1.06$ $35.60$ $0.975$ $0.040$ $3.8$
+Complete model $2.32$ $24.92$ $0.906$ $0.122$ $\underline{0.94}$ $\underline{36.07}$ $0.977$ $0.038$ $6.6$
+w/o $L_{\text{over}}$ $\underline{2.11}$ $\underline{25.17}$ $\underline{0.914}$ $\underline{0.118}$ $1.34$ $36.04$ $\underline{0.977}$ $\underline{0.037}$ $7.6$
+w/o $L_{\text{opa}}$ $5.72$ $21.12$ $0.853$ $0.174$ $1.06$ $35.60$ $0.975$ $0.040$ $3.8$
 w/o $L_{\text{cov}}$ $3.21$ $22.96$ $0.896$ $0.134$ $1.34$ $36.00$ $0.976$ $0.038$ $8.53$
 w/o Adaptive $3.16$ $22.74$ $0.880$ $0.141$ $1.05$ $35.74$ $0.974$ $0.040$ $6.7$
-w/o $L_{\text{par}}$ 1.91 24.47 0.898 0.119 0.93 36.18 0.977 0.037 $10.1$
-IV-C3 Real-life Data
+w/o $L_{\text{par}}$ 1.91 25.48 0.918 0.115 0.93 36.18 0.977 0.037 $10.1$
+Table 5: Ablation studies on the ShapeNet [6]. We report Chamfer Distance, rendering metrics, and the number of primitives (#P).
+4.3.2 Real-life Data
 To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from BlenderMVS dataset [45] and a self-captured dataset.
 As shown in Fig.
-5, our approach can robustly produce both realistic appearances and meaningful 3D decompositions across a variety of data types.
-Additionally, the distinct advantages of block-level and point-level reconstructions are clearly observable.
-For instance, in the sixth column, due to limited training views, point-level reconstructions based on 2DGS tend to generate artifacts, whereas block-level optimization produces accurate and smooth surfaces.
-Conversely, point-level reconstruction demonstrates superior capability in capturing intricate geometric details.
-Figure 7: Impact of Gaussian scale regularization (SR).
-This figure demonstrates the effect of Gaussian scale regularization.
-As shown by the typically degraded mesh reconstruction from 2DGS [17], scale regularization effectively fixes the holes in the ShapeNet [57] dataset.
-Figure 8: Impact of the initial number of primitives.
-We qualitatively demonstrate the impact of the initial number of primitives on the granularity of decomposition.
-A larger number of initial primitives results in a finer-grained decomposition, whereas a smaller number leads to a coarser decomposition.
+4, our approach can robustly produce both realistic appearances and meaningful 3D decompositions across a variety of data types.
+More results are provided in the supplementary material.
+Figure 7: Impact of Gaussian scale regularization (SR). The degraded mesh produced by 2DGS [16] is effectively improved.
+Figure 8: Impact of the initial number of primitives ($M$). Increasing $M$ yields a finer-grained decomposition, while decreasing it produces a coarser decomposition.
 Figure 9: Applications.
-Part-Aware Editing (top): After optimization, we can easily edit the scene by moving, adding, removing, or scaling specific parts.
-3D Content Generation (bottom): By combining parts of different objects, we can effortlessly create new and meaningful 3D content.
-IV-D Ablations
-TABLE VI: Effect of the parsimony loss weight ($\lambda_{\text{par}}$) on DTU [55].
-We report Chamfer Distance (CD), rendering metrics, and the number of primitives (#P) of block-level reconstruction and point-level reconstruction.
+Part-Aware Editing (top): After optimization, we can easily edit the scene by adding, scaling, or moving specific parts.
+3D Content Generation (middle): By combining parts of different objects, we can create new 3D content.
+Physics simulations (bottom): The representation enables us to perform physics-based simulations easily.
+4.3.3 Compared to SAM-based Methods
+We also conduct comparisons to SAM-based methods for this task, as shown in Fig.
+6.
+Despite impressive advances in 3D segmentation and editing [5, 19, 46, 47] by distilling 2D information, achieving finer-grained semantic disentanglement remains challenging due to indistinct textures and severe cross-view 2D inconsistencies.
+Lifting 2D segmentation features to 3D segmentation will also lead to feature misalignment, leading to visual artifacts.
+In contrast, our method segments directly from 3D space, enabling more fine-grained and clearly defined component decomposition.
+4.4 Ablations
 Block-level Point-level
@@
 $\lambda_{\text{par}}$ = 0.1 $6.46$ $16.10$ $0.721$ $0.255$ $1.08$ $34.30$ $0.987$ $0.018$ $2.7$
-$\lambda_{\text{par}}$ = 0.01 (default) $4.19$ $19.84$ $0.820$ $0.189$ $1.05$ $35.04$ $0.988$ $0.015$ $5.9$
+$\lambda_{\text{par}}$ = 0.01 (*) $4.19$ $19.84$ $0.820$ $0.189$ $1.05$ $35.04$ $0.988$ $0.015$ $5.9$
 $\lambda_{\text{par}}$ = 0.001 $4.01$ $20.18$ $0.835$ $0.180$ $1.03$ $35.04$ $0.988$ $0.015$ $7.1$
-TABLE VII: Effect of the initial number of primitives ($M$) on DTU [55].
-We report Chamfer Distance (CD), rendering metrics, and the number of primitives (#P) of block-level reconstruction and point-level reconstruction.
-Block-level Point-level
-[…]
 $M$= 4 $4.97$ $19.06$ $0.794$ $0.213$ $1.06$ $34.93$ $0.987$ $0.016$ $5.0$
-$M$= 8 (default) $4.19$ $19.84$ $0.820$ $0.189$ $1.05$ $35.04$ $0.988$ $0.015$ $5.9$
+$M$= 8 (*) $4.19$ $19.84$ $0.820$ $0.189$ $1.05$ $35.04$ $0.988$ $0.015$ $5.9$
 $M$= 16 $3.99$ $20.87$ $0.83$ $0.176$ $1.07$ $35.05$ $0.988$ $0.015$ $8.3$
-In this section, we first verify and analyze the critical strategies proposed by our algorithm: overlapping loss, opacity loss, coverage loss, adaptive number of primitives, and parsimony loss.
-By removing each strategy one at a time, we report the averaged quantitative performance of reconstruction on 15 instances of chair category in Tab.
-V and provide visual comparisons in Fig.
-6.
-As shown in Tab.
-V, VI, and VII, the point-level reconstruction accuracy and rendering quality do not differ significantly.
-This is because the refinement stage primarily assists in recovering the details of geometry rather than part decomposition.
-Therefore, in the subsequent analysis, we will focus on the block-level reconstruction results.
-Removing the overlapping loss improves reconstruction quality and reduces decomposition semantics while retaining more parts.
-Intuitively, more superquadrics will better fit a shape, but overlaps diminish the meaningfulness of decomposition, as shown in the first column of Fig.
-6.
-Removing opacity loss drastically affects the number of primitives, reducing them to an average of only 3.8.
-Simultaneously, the CD increases by approximately twofold.
-This demonstrates the importance of opacity loss in controlling the existence of primitives and ultimately impacting reconstruction accuracy.
-In the visual example, the reconstructed chair is missing its legs and armrests.
-Coverage loss significantly involves reconstruction errors as the CD increases by 38%, ensuring the primitives stay in the correct position to fit the target best.
-As shown in the third column of Fig.
-6, there is space in the chair seat that primitives do not cover without coverage loss.
-The adaptive number of primitives strategy fills in missing parts and improves the reconstruction metric by 36%.
-Omitting parsimony loss results in the lowest CD errors and best rendering metrics but produces 0.5 times more primitives than the full model, indicating redundancy and over-decomposition.
+Table 6: Effect of parsimony weight ($\lambda_{\text{par}}$) and initial primitives count ($M$) on DTU [17].
+We report Chamfer distance, rendering metrics, and primitives count (#P).
+* denotes the default setting.
+In this section, we first analyze the design choices by isolating each strategy to assess its impact.
+We report the averaged quantitative performance on 15 instances of chair category in Tab.
+5 and provide visual comparisons in Fig.
+5.
+Removing overlapping loss enhances reconstruction quality but reduces semantic clarity.
+Omitting opacity loss drastically affects the number of primitives and doubles the CD, highlighting its role in controlling primitive presence and reconstruction accuracy.
+Coverage loss is crucial in reconstruction accuracy, ensuring primitives align correctly with the target.
+The adaptive primitive strategy fills in missing parts and improves the reconstruction metric.
+Without parsimony loss, CD and rendering metrics are optimal but result in redundancy and over-decomposition.
 We also validate the effectiveness of Gaussian scale regularization in Fig.
 7.
-We find that 2DGS [17] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF.
-The holes occur since large-scale Gaussians can’t support consistent depth rendering from different views.
-Scale regularization effectively fixes these holes by preventing the formation of large-scale Gaussians.
-Lastly, we analyze the impact of two key hyperparameters of our method on decomposition granularity: the parsimony loss weight $\lambda_{\text{par}}$ and the initial number of primitives $M$.
-As shown in Tab.VI, $\lambda_{\text{par}}$ directly influences the number of primitives found (#P).
-Stronger regularization results in fewer primitives, while weaker regularization has the opposite effect.
-However, having too few primitives severely impacts reconstruction accuracy and rendering quality.
-In Tab.VII, it can be observed that as $M$ increases, both reconstruction and view synthesis performance slightly improve.
-Simultaneously, $M$ positively correlates with #P.
-This indicates that by adjusting the values of $M$ and $\lambda_{\text{par}}$, our approach can control the granularity of object or scene decomposition.
+2DGS [16] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF.
+Holes occur since large-scale Gaussians can’t support consistent depth rendering from different views.
+Scale regularization effectively addresses this issue by suppressing large-scale Gaussians.
+Lastly, in Tab.
+6, we analyze the influence of two key hyperparameters on decomposition granularity: the parsimony loss weight $\lambda_{\text{par}}$ and the initial primitive count $M$.
+Stronger parsimony regularization reduces the number of primitives, while weaker regularization increases it.
+As $M$ rises, both reconstruction and view synthesis performance slightly improve.
+This demonstrates that adjusting $M$ and $\lambda_{\text{par}}$ enables our method to effectively control the granularity of object or scene decomposition.
 Fig.
-8 visually represents this impact.
-IV-E Applications
+8 visually illustrates this impact.
+4.5 Applications
 Fig.
-9 illustrates two applications of our method that original 3DGS and NeRF-based approaches do not support.
-Firstly, after optimization, we obtain the part decomposition, making it easy to edit specific parts of the object or scene, e.g., adding, moving, removing, or scaling sizes.
+9 illustrates three applications of our method that original 3DGS and NeRF-based approaches do not support.
+Firstly, after optimization, we obtain the part decomposition, facilitating easy editing of specific object or scene components, e.g., adding, moving, removing, or scaling.
 Secondly, by combining parts of different objects, our method enables the creation of new high-quality 3D content.
-By replacing parts, we do not need to adjust the position and part connections manually.
-For man-made objects, the proposed method supports 3D shape generation within the same categories nicely.
-V Conclusion
+Additionally, our method enables easy manipulation of objects within a scene, supporting physics-based simulations.
+5 Conclusion
 We introduce a hybrid representation of superquadrics and 2D Gaussians, to learn 3D scenes in a part-aware representation.

### v4

@@
-PartGS: Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics
+Self-supervised Learning of Hybrid Part-aware 3D Representations of 2D Gaussians and Superquadrics
 Abstract
-Low-level 3D representations, such as point clouds, meshes, NeRFs, and 3D Gaussians, are commonly used to represent 3D objects or scenes.
-However, human perception typically understands 3D objects at a higher level as a composition of parts or structures rather than points or voxels.
-Representing 3D objects or scenes as semantic parts can benefit further understanding and applications.
-In this paper, we introduce PartGS, part-aware 3D reconstruction by a hybrid representation of 2D Gaussians and Superquadrics, which parses objects or scenes into semantic parts, digging 3D structural clues from multi-view image inputs.
-Accurate structured geometry reconstruction and high-quality rendering are achieved at the same time.
-Our method simultaneously optimizes superquadric meshes and Gaussians by coupling their parameters within our hybrid representation.
-On one hand, this hybrid representation inherits the advantage of superquadrics to represent different shape primitives, supporting flexible part decomposition of scenes.
-On the other hand, 2D Gaussians capture complex texture and geometry details, ensuring high-quality appearance and geometry reconstruction.
-Our method is fully unsupervised and outperforms existing state-of-the-art approaches in extensive experiments on DTU, ShapeNet, and real-life datasets.
+Low-level 3D representations, such as point clouds, meshes, NeRFs and 3D Gaussians, are commonly used for modeling 3D objects and scenes.
+However, cognitive studies indicate that human perception operates at higher levels and interprets 3D environments by decomposing them into meaningful structural parts, rather than low-level elements like points or voxels.
+Structured geometric decomposition enhances scene interpretability and facilitates downstream tasks requiring component-level manipulation.
+In this work, we introduce PartGS, a self-supervised part-aware reconstruction framework that integrates 2D Gaussians and superquadrics to parse objects and scenes into an interpretable decomposition, leveraging multi-view image inputs to uncover 3D structural information.
+Our method jointly optimizes superquadric meshes and Gaussians by coupling their parameters within a hybrid representation.
+On one hand, superquadrics enable the representation of a wide range of shape primitives, facilitating flexible and meaningful decompositions.
+On the other hand, 2D Gaussians capture detailed texture and geometric details, ensuring high-fidelity appearance and geometry reconstruction.
+Operating in a self-supervised manner, our approach demonstrates superior performance compared to state-of-the-art methods across extensive experiments on the DTU, ShapeNet, and real-world datasets.
+footnotetext: Corresponding authors
 1 Introduction
-Figure 1: We propose a pipeline to reconstruct the scene in a part-aware manner.
-Prior works [22, 27] model the scene using primitives, while the proposed method can further model precise geometry details and textures.
+Figure 1: Prior works [32, 36] model the scene using primitives, while the proposed method can further model precise geometry details and textures.
 Here, EMS [32] takes point cloud inputs and reconstructs non-textured primitives.
-3D reconstruction from multi-view images is a long-studying problem in 3D vision and computer graphics.
-Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes, which differ from human perception.
-Humans understand 3D scenes as different semantic parts [26].
-For instance, when observing a scene, we mentally construct high-level structural information, such as scene graphs, rather than point clouds or detailed geometry.
-Thus, we propose to learn part-aware scene reconstruction to decompose scenes into different individual semantic shapes or parts, facilitating tasks including physical simulation, manipulation/editing, 3D content generation, etc.
-There are prior methods [38, 51, 2, 28, 20, 44, 22, 23, 36, 42, 48] working on this topic, while they mainly learn the goal with 3D supervisions and cannot preserve accurate geometries, which leaves inconvenience to apply in realistic scenarios.
-While NeRF [25] has demonstrated great potential in reconstructing textured 3D scenes from multi-view images, some approaches [37, 49] try to learn part-aware objects via NeRFs.
-In particular, Part-NeRF [37] represents an object as multiple neural radiance fields.
-However, the complex composition of multiple implicit fields increases learning difficulty, resulting in subpar rendering quality and insufficient semantic decomposition.
-Another state-of-the-art approach DBW [27] decomposes scenes into blocks represented by superquadric primitives [2], optimizing their parameters and UV texture maps by minimizing rendering loss.
-Although DBW effectively decomposes 3D scenes into reasonable parts, learning accurate part-aware geometry and appearance still proves challenging as shown in Fig.
+3D reconstruction from multi-view images is a long-standing challenge in 3D vision and graphics [55, 17, 12].
+Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes.
+However, humans tend to understand 3D scenes as reasonable parts [35].
+For instance, when observing a scene, we naturally construct high-level structural information, such as scene graphs, instead of focusing on low-level details like point clouds or voxels.
+Motivated by it, we propose a part-aware reconstruction framework that decomposes objects or scenes into meaningful shapes or parts, facilitating tasks such as physical simulation, editing, content generation and understanding [18, 28, 58, 53, 54].
+Several prior works [72, 37, 29, 32, 33, 56] have explored part-aware reconstruction or 3D decomposition.
+However, most of them rely heavily on 3D supervision and often struggle to retain fine-grained geometric details, limiting their practicality in real-world scenarios.
+Recent advances [48, 68] have focused on extending the neural radiance field framework for part-aware reconstruction, building upon its remarkable success in 3D reconstruction from multi-view images.
+For example, PartNeRF [48] models objects using multiple neural radiance fields.
+Yet, the intricate composition of implicit fields complicates the learning process, leading to suboptimal rendering quality and inefficient decomposition.
+Recently, DBW [36] proposes a novel approach that decomposes scenes into block-based representations using superquadric primitives [2], optimizing both their parameters and UV texture maps through rendering loss minimization.
+While this method demonstrates effective scene decomposition into coherent components, it faces challenges in achieving accurate part-aware reconstruction of both geometry and appearance, as illustrated in Fig.
 1.
-To tackle the problem, we propose PartGS, a part-aware hybrid representation that integrates 2D Gaussians [16] and superquadrics [2] to reconstruct scenes that are both well-textured and geometrically accurate.
+To address these limitations, we propose PartGS, a self-supervised part-aware hybrid representation that integrates 2D Gaussians [21] and superquadrics [2], to achieve both high-quality texture reconstruction and geometrically accurate part decomposition.
 There are previous approaches [20, 14, 50] combining mesh reconstruction and Gaussian splatting, where they first reconstruct the mesh and then attach Gaussians to it.
-Their purpose is to get both representations, while our goal is to get part segmentation and reconstruction at the same time.
-Our method directly learns superquadric mesh and Gaussians simultaneously by coupling parameters between them while constraining each part to be represented as a single superquadric.
-The rendering capability of coupled Gaussians drives the optimization of hybrid representation, enabling the use of their convexity to assist in the decomposition of shapes.
-The unsupervised training of part-aware reconstruction is built upon the assumption that each functional part of most objects should be a basic shape and can be represented by a superquadric.
-On the other hand, Gaussian splatting, renowned for its superior rendering capabilities and efficient training[18, 11, 24], is incorporated to capture intricate texture details, speeding up reconstruction and improving rendering quality significantly.
-In the hybrid representation, 2D Gaussians are attached to the surface of superquadrics, with their pose and shape parameters coupled to the corresponding superquadric mesh.
-Compared to the vanilla Gaussian [18] that fills up the whole occupancy with individual-parameter 3D Gaussians, our method is more compact and efficient.
-The optimized parameters include global ones controlling the overall shape, pose, and opacity of hybrid blocks, and local parameters for spherical harmonic coefficients of the Gaussians.
-During training, the parameters are optimized by rendering loss without additional supervision.
-We observe that image rendering loss tends to settle into local minima during the optimization of superquadrics.
-To address this, we design several regularizers to enforce global consistency between the 3D representation and the input 2D information.
-This allows us to segment parts in a fully unsupervised manner, achieving block-level reconstruction.
-At last, to model the accurate geometry, we conduct a point-level refinement, freeing 2D Gaussians to get out of the surface, thereby improving the modeling of irregular shapes.
-Contributions are summarized as follows:
-  • A novel hybrid representation is introduced to learn part-aware textured 3D scenes.
-The hybrid representation combines the benefits of both superquadrics and Gaussian splitting, enabling accurate part reconstruction and efficient high-quality rendering.
-  • An end-to-end unsupervised pipeline is proposed for block-level and point-level part-aware reconstruction, with superquadric meshes and 2D Gaussians optimized simultaneously by novel regularizers.
-  • Extensive experiments on public datasets demonstrate the superiority of the proposed method in part-aware reconstruction, enabling applications of part augmentation and physical interactions.
-Figure 2: Overview of PartGS.
-PartGS takes multi-view images as input to learn a parametric hybrid representation of superquadrics and 2D Gaussians.
-The representation initializes from random superquadrics and is gradually optimized during training to get a block-level reconstruction.
-Then, we free the constraints of Gaussians to model detailed geometry to get a point-level reconstruction.
-The last column demonstrates one application of PartGS: physics simulations, where one part is picked up from the scene and thrown forward.
-Please refer to the supplementary materials for more video demos.
-2 Related Works
-2.1 Primitive-based 3D Representation
-Learning structured shape representations aims to decompose objects or scenes into semantically coherent geometric primitives, providing a crucial pathway to shape understanding and generation [43, 50, 14, 9, 41, 35].
-The earliest classical approaches, dating back to the last century, such as Blocks World [32] and Generalized Cylinders [3] emphasize compactness as the central goal.
-A substantial amount of prior work has focused on processing 3D data inputs, such as point clouds, meshes, and voxel grids.
-These methods decompose the 3D data into various types of primitive ensembles, including cuboids [38, 31], superquadrics [29, 30, 44, 22], and convex shapes [8, 10].
-MonteBoxFinder [31], uses a RANSAC-based [12] with cuboids to represent point clouds, combined with Monte Carlo Tree Search for optimal selection.
-Alternatively, EMS [22] employs superquadrics to represent geometric primitives and proposes a probability-based approach to recover primitives.
-These methods rely on perfect 3D data, which is rarely available due to incompleteness or noise.
-Few studies [37, 27, 49, 1] have attempted to create structure-aware 3D representations directly from images.
-PartNeRF [37] uses ellipses with neural radiance fields, but this spatial-based decomposition limits its ability to capture meaningful structural information.
+In contrast to their primary focus on acquiring the representations, the proposed approach is to achieve part-aware decomposition reconstruction.
+Our method establishes a coupled optimization framework that simultaneously learns superquadric meshes and Gaussians through parameter sharing, with each part constrained to a single superquadric.
+The differentiable rendering of Gaussians drives this hybrid representation, leveraging the inherent convexity of superquadrics for qualitative 3D decomposition.
+The self-supervised training of part-aware reconstruction is built upon the assumption that each part of most objects should be a basic shape and can be represented by a superquadric.
+Meanwhile, Gaussian splatting, renowned for its superior rendering quality and efficient training, is incorporated to capture intricate texture details, speeding up reconstruction and improving rendering quality.
+In the hybrid representation, 2D Gaussians are distributed on superquadric surfaces to form structured blocks.
+The pose and shape of Gaussians within each block are determined by their corresponding superquadrics rather than being optimized independently.
+The parameter space encompasses global controls for superquadric properties (shape, pose, and opacity) and local spherical harmonic coefficients for individual Gaussians.
+Compared to standard Gaussian splatting [25, 21], which populates the occupancy volume with independently parameterized Gaussians, the coupled representation is more compact and efficient.
+During training, the parameters are optimized through rendering loss without additional supervision.
+However, we notice that image rendering loss often leads to local minima in superquadric shape optimization.
+To tackle this, we introduce several regularizers to maintain global consistency between the 3D representation and the input 2D information.
+This strategic implementation allows us to segment parts fully self-supervised, achieving block-level reconstruction.
+Finally, to better capture irregular shapes, we implement a point-level refinement step that frees 2D Gaussians to deviate from the surface, thereby enhancing geometric fidelity.
+Extensive experiments show that PartGS, at block-level and point-level, achieves 33.3% and 75.9% improvements in reconstruction accuracy, 3.18 and 16.13 increases in PSNR, and 4X and 3X speedups compared to the state-of-the-art baseline.
+Our contributions are summarized as follows:
+  • We introduce a novel hybrid representation for part-aware 3D reconstruction, combining the strengths of superquadrics and Gaussian splatting to achieve reasonable decomposition and high-quality rendering.
+  • We propose several novel regularizers to enforce consistency between 3D decomposition and 2D observations, enabling self-supervised part decomposition.
+  • Compared to prior works, the method takes one step forward, by reconstructing both the block-level and point-level part-aware reconstructions, preserving both part segmentation and reconstruction precision.
+Figure 2: Overview of our pipeline.
+PartGS takes multi-view images to learn a parametric hybrid representation of superquadrics and 2D Gaussians.
+It initializes from random superquadrics and is gradually optimized during training to obtain a block-level reconstruction.
+Then, we free the constraints of Gaussians to model detailed geometry, achieving point-level reconstruction.
+2 Related Work
+2.1 Shape Decomposition and Abstraction
+Structured shape representation learning decomposes objects or scenes into coherent geometric primitives, facilitating shape understanding and generation [59, 19, 15, 9, 47].
+Early approaches like Blocks World [43] and Generalized Cylinders [3] emphasized compact representations.
+Modern methods typically process 3D inputs (point clouds, meshes, voxels) by decomposing them into primitive ensembles, including cuboids [49, 42], superquadrics [38, 39, 60, 32], and convex shapes [8, 10].
+For instance, MonteBoxFinder [42] integrates clustering [11, 51, 62], cuboid representations, and Monte Carlo Tree Search, while EMS [32] adopts a probabilistic approach for superquadric recovery.
+However, these methods are limited to coarse shape representations.
+Recent advances in shape abstraction [13, 40, 22, 46, 31] enable more detailed shape representations through flexible primitive deformation.
+Some studies [48, 36, 68, 1, 16, 70] attempt to create structure-aware 3D representations directly from images.
+PartNeRF [48] introduces ellipsoid representations within NeRFs, its reliance on multiple implicit fields results in inefficient 3D decomposition.
 ISCO [1] and DBW [36] use 3D superquadrics for shape decomposition, which enables more meaningful structural separation.
@@
 DPA-Net [68] has advanced 3D shape abstraction from sparse views but generates redundant parts and struggles with realistic texture rendering.
-Our approach introduces a hybrid representation that learns part-aware 3D scenes while ensuring accurate structural reconstruction and efficient, high-quality rendering.
+A concurrent work, GaussianBlock [24], employs SAM [27] to guide superquadric splitting and fusion for 3D decomposition, yet its computational efficiency remains limited, typically requiring several hours per scene.
+In contrast, our approach accomplishes self-supervised [61, 52] part-aware scene reconstruction through an efficient hybrid representation that simultaneously maintains geometric fidelity and photorealistic rendering quality.
 2.2 Mesh-based Gaussian Splatting
-Gaussian splatting(GS) [18] has been rapidly adopted in multiple fields due to its remarkable rendering capability.
+Gaussian splitting (GS) [25] has been rapidly adopted in multiple fields due to its remarkable rendering capability.
 Several studies [20, 7, 14, 50] attempt to align Gaussians with mesh surfaces for easier editing and animation.
@@
 3 Method
-Given a set of calibrated multi-view images, our goal is to learn part-aware 3D geometry and appearance.
-A part-aware reconstruction is often represented as a composition of multiple basic shapes (geometric primitives).
-We propose a primitive-based hybrid representation in Sec.
-3.1.
-In Sec.
-3.2, we demonstrate how this hybrid representation is optimized to fit primitives as parts (block-level).
-In order to further detail the geometry of each part, the parameters of Gaussians are further refined to get a point-level reconstruction.
-3.1 Hybrid Representation
+Given a set of calibrated multi-view images and foreground masks, we aim to learn part-aware 3D representations, a meaningful decomposition of both geometry and appearance.
+Our approach adopts a two-stage optimization strategy: first, decomposing the object into basic shapes using a mixture of Gaussians and superquadrics at the block-level, followed by refining the decomposition at the point-level to achieve precise geometry.
+In Sec 3.1, we parameterize the hybrid representation.
+Sec.
+3.2 then elaborates on leveraging this representation, enhanced by novel regularizers and an adaptive control strategy, to achieve self-supervised block-level decomposition.
+Finally, Sec.
+3.3 presents the process for obtaining detailed part-aware results.
+3.1 Parametrizing the Hybrid Representation
 As shown in Fig.
-2, the proposed hybrid representation combines superquadrics with 2D Gaussians in a compact form.
-To leverage the strengths of both, we attach Gaussians to the surface of superquadric meshes.
-This representation retains the superquadric’s ability to effectively parse a 3D scene into distinct parts.
-Meanwhile, spherical harmonics of Gaussians enable complex texture rendering via Gaussian splatting, addressing the texture learning limitations of previous work [22, 31, 27, 48].
-Sharing pose parameters between superquadrics and 2D Gaussians further enhances the representation’s efficiency.
-The proposed hybrid model is parametric, controlled by both primitive and Gaussian parameters, which are optimized simultaneously.
-Considering a 3D scene $\mathcal{S}$, we parse it into multiple superquadrics, i.e.
-blocks.
-Each block is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$.
-Here, $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks.
-The hybrid blocks are defined using manually designed parameters that control pose, opacity, scale, shape, and texture.
+2, to leverage the strengths of both, we attach Gaussians to the surface of superquadric meshes.
+This representation retains the superquadric’s ability to parse a 3D scene into distinct parts.
+Meanwhile, spherical harmonics of Gaussians enable complex texture rendering via Gaussian splatting, addressing the texture learning limitations of prior work [32, 42, 36, 67].
+Sharing pose parameters between superquadrics and 2D Gaussians further improves the representation’s efficiency.
+The parametric representation is controlled by both primitive and Gaussian parameters, which are optimized simultaneously.
+Given a 3D scene $\mathcal{S}$, the proposed method decomposes it into multiple hybrid blocks, each consisting of a superquadric with associated Gaussians.
+Each scene is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$, where $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks.
+Blocks are defined using manually designed parameters that control pose, opacity, scale, shape, and texture.
 These parameters are optimized via differentiable rendering to parse the target scene.
@@
 For each hybrid block $\mathcal{B}_{i}$, its geometry is controlled by superquadric parameters [2].
-More concretely, there are two shape parameters $\epsilon_{1},\epsilon_{2}$ that define the superquadric mesh, along with three scale parameters $s_{1},s_{2},s_{3}$, which scale the 3D axes.
-Similar to the positioning of vertexes in an icosphere, the 3D position of a superquadric vertex can be obtained by:
+Specifically, there are two shape parameters $\epsilon_{1},\epsilon_{2}$ that define its shape, along with three scale parameters $s_{1},s_{2},s_{3}$, which scale the 3D axes.
+Analogous to icosphere vertex positioning, superquadric vertex coordinates are computed by:
 $\footnotesize\mathbf{v}=[s_{1}\cos^{\epsilon_{1}}(\theta)\cos^{\epsilon_{2}}(\varphi);s_{2}\sin^{\epsilon_{1}}(\theta);s_{3}\cos^{\epsilon_{1}}(\theta)\sin^{\epsilon_{2}}(\varphi)],$ (1)
-where $\theta$ and $\varphi$ represent the azimuth and elevation defined in the spherical coordinate, respectively.
-The shape and scale parameters are optimized during training, thus controlling the block deformation to learn part-aware geometry.
+where $\theta$ and $\varphi$ represent the azimuth and elevation defined in the spherical coordinate. The shape and scale parameters govern block deformation to learn part-aware geometry.
 Pose Parameters.
@@
 1 is transformed from the local coordinate to world space as: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $j$ indexes the vertices.
-In prior methods such as SuGaR [15] and GaMeS [39], Gaussians are placed on the reconstructed mesh, and they have individual poses and shapes for better appearance modeling.
-Differently, in our method, the role of Gaussians is to construct a differentiable rendering that bridges images and superquadrics.
+Previous approaches, including SuGaR [20] and GaMeS [50], position Gaussians on reconstructed meshes with independent pose and shape parameters to enhance appearance modeling.
+In contrast, our method employs Gaussians to construct a differentiable rendering that bridges images and superquadrics.
 To achieve this, we couple the parameters of Gaussians with superquadrics.
-Given a posed superquadric, centers of Gaussians are determined by uniformly sampling on triangular faces, and poses of Gaussians are determined by face vertices.
-Inspired by GaMeS [39], we assign the rotation matrix $\mathrm{R}_{v}$ and scaling $\mathrm{S}_{v}$ of Gaussians based on vertices’ positions.
-For a given triangular face $V$ and its vertices $v_{1}$, $v_{2}$, $v_{3}\in\mathbb{R}^{3}$, we define orthonormal vectors $r_{1}$, $r_{2}$, $r_{3}\in\mathbb{R}^{3}$ to construct the rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$.
-The vector $r_{1}$ is aligned with the normal vector of $V$ to ensure Gaussians align with the triangular surface.
-The vector $r_{2}$ is defined as the vector from the center $m=\text{mean}(v_{1},v_{2},v_{3})$ of the triangle to the first vertex $v_{1}$:
-$r_{2}=\frac{v_{1}-m}{||v_{1}-m||}.$ (2)
-The last rotation vector $r_{3}$ is obtained by orthonormalizing the vector from the center to the second vertex concerning the existing $r_{1}$ and $r_{2}$ :
+For a posed superquadric, Gaussian centers are uniformly sampled on triangular faces, with their poses determined by face vertices.
+Following GaMeS [50], each Gaussian’s rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$ and scaling $\mathrm{S}_{v}$ are computed from vertex positions.
+Given a triangular face $V$ with vertices $v_{1},v_{2},v_{3}\in\mathbb{R}^{3}$, orthonormal vectors are constructed such that $r_{1}$ aligns with the face normal, $r_{2}$ points from the centroid $m=\text{mean}(v_{1},v_{2},v_{3})$ to $v_{1}$.
+$r_{3}$ is obtained by orthonormalizing the vector from the center to the second vertex concerning the existing $r_{1}$ and $r_{2}$ :
 $r_{3}=\frac{{ort}(v_{2}-m;r_{1},r_{2})}{||{ort}(v_{2}-m;r_{1},r_{2}||},$ (2)
@@
 We place a fixed number of Gaussians on each triangular face.
-These designs eliminate the need to learn the geometry of Gaussians, focusing solely on their texture, thereby enhancing the efficiency of the representation.
+These designs eliminate the need to learn the geometry of Gaussians, thereby enhancing the efficiency of the representation.
 Opacity Parameters.
 We define the total number of hybrid blocks as $M$.
-However, in practice, a typical scene does not contain exactly $M$ blocks.
+However, a typical scene does not contain exactly $M$ blocks.
 Therefore, only the meaningful blocks are retained.
 To achieve this, we introduce a learnable parameter $\tau_{i}$ to represent each block’s opacity.
-During the optimization process, only blocks with $\tau_{i}$ greater than a certain threshold are retained.
-Note that Gaussians on the same block share the same $\tau$ for opacity in rasterization.
+During the optimization, only blocks with $\tau_{i}$ greater than a certain threshold are retained.
+Note that Gaussians within the same block share the same $\tau$ for opacity in rasterization.
 Texture Parameters.
@@
 Spherical harmonics of each Gaussian control the texture and are optimized for rendering view-dependent images [25].
-3.2 Optimizing the Hybrid Representation
+3.2 Block-level Decomposition
 This section describes the optimization of the hybrid representation.
-We found that minimizing the rendering loss across multi-view images alone led to instability in positioning hybrid blocks (primitives) accurately.
-Therefore, we introduce several regularizations and perform two optimization stages: block-level and point-level optimization.
-3.2.1 Block-level Optimization
+We observed that minimizing the rendering loss across multi-view images alone led to instability in positioning hybrid blocks.
+Therefore, several regularization terms are introduced to optimize the composition for maximal image formation consistency.
+3.2.1 Optimization
 In addition to the image rendering loss, we employ regularizers consisting of coverage, overlap, parsimony, and opacity entropy.
 The coverage loss encourages hybrid blocks to cover only meaningful regions; the overlap loss prevents blocks from overlapping; the parsimony loss regularizes the number of existing blocks; and the opacity entropy encourages block opacities close to binary.
-Rendering Loss. Based on the rendering process of Gaussian spaltting [18], the rendering loss is calculated by associating $L_{1}$ with a D-SSIM term[18] :
+Rendering Loss. The rendering loss is based on 3DGS [25], where it combines an $L_{1}$ term with a D-SSIM term:
 ${\mathcal{L}}_{\text{ren}}=(1-\lambda)\ L_{1}+\lambda L_{\text{D-SSIM}}.$ (4)
 Coverage Loss.
-This loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending beyond the object’s boundaries.
-To determine the 3D occupancy field based on the blocks, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=1-\Psi_{i}(x)$.
+The coverage loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending beyond its boundaries.
+To determine the 3D occupancy field based on the blocks, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=\Psi_{i}(x)-1$.
 Here, $\Psi_{i}$ denotes the superquadric inside-outside function [2] associated with the block $i$.
-Consequently, $D_{i}(x)\geq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)<0$ if $x$ is outside the block.
+Consequently, $D_{i}(x)\leq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)>0$ if $x$ is outside the block.
 Inspired by NeRF [34], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses.
@@
 ${\mathcal{L}}_{\text{cov}}({\mathcal{R}})=\sum_{r\in{\mathcal{R}}}l_{r}L_{\text{cross}}(r)+(1-l_{r})L_{\text{non-cross}}(r).$ (5)
-Here, $L_{\text{cross}}$ encourages rays to intersect with blocks, while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks. They are defined as:
-$L_{\text{cross}}(r)=\text{ReLU}\Big{(}\min_{i\in\mathcal{M}}\min_{x_{j}^{r}\in{\mathcal{X}}_{r}}D_{j}(x_{j}^{r})\Big{)},$ (7)
+Here, $L_{\text{cross}}$ encourages rays to intersect with blocks, while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks:
+$L_{\text{cross}}(r)=\text{ReLU}\Big{(}\min_{i\in\mathcal{M}}\min_{x_{j}^{r}\in{\mathcal{X}}_{r}}D_{i}(x_{j}^{r})\Big{)},$ (6)
 $L_{\text{non-cross}}(r)=\text{ReLU}\Big{(}\max_{i\in\mathcal{M}}\max_{x_{j}^{r}\in{\mathcal{X}}_{r}}-D_{i}(x_{j}^{r})\Big{)},$ (7)
-where $x_{j}^{r}$ means the $j$-th sampled points along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$.
-Intuitively, this means that there is at least one in sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not lie within any block.
+where $x_{j}^{r}$ means the $j$-th sampled point along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$.
+Intuitively, this implies that at least one sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not belong to any block.
 Overlap Loss.
-We add a regularization term to encourage individual blocks to remain non-overlapping.
-Given the difficulty of directly calculating block overlap, we use a Monte Carlo method similar to  [27].
-Practically, we sample multiple 3D points in space and penalize each point that lies inside more than $k$ blocks.
-Based on the definition of the superquadric inside-outside function, we devise a soft occupancy function:
-$\mathcal{O}_{i}^{x}=\tau_{i}\text{ sigmoid}(\frac{1-\Psi_{i}(x)}{\gamma}),$ (9)
-where $\gamma$ is a temperature parameter. Therefore, given a set of $N$ sampled 3D points $\Omega$, the overlap loss is expressed as:
+We introduce a regularization term to prevent overlap between individual blocks.
+Given the difficulty of directly calculating block overlap, we adopt a Monte Carlo method similar to  [36].
+Specifically, we sample multiple 3D points in space and penalize those that lie inside more than $k$ blocks.
+Based on the superquadric inside-outside function, a soft occupancy function is defined as:
+$\mathcal{O}_{i}^{x}=\tau_{i}\text{ sigmoid}(\frac{-D_{i}(x)}{\gamma}),$ (8)
+where $\gamma$ is a temperature parameter. Thus, for a set of $N$ sampled 3D points $\Omega$, the overlap loss is expressed as:
 ${\mathcal{L}}_{\text{over}}=\frac{1}{N}\sum_{x\in\Omega}\text{ReLU}\left(\sum_{i\in\mathcal{M}}\mathcal{O}_{i}^{x}-k\right).$ (9)
-Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we add a loss term that regularizes the block opacity ($\tau$). This loss is defined as:
+Parsimony Loss.
+To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we introduce a regularization term that penalizes block opacity ($\tau$).
+This loss is defined as:
 ${\mathcal{L}}_{\text{par}}=\frac{1}{M}\sum_{i\in\mathcal{M}}\sqrt{\tau_{i}}.$ (10)
 Opacity Entropy Loss.
-Naturally, during optimization, the opacity of the block inside the object region tends to approach 1, while the opacity of the block outside the object region tends to approach 0.
-To facilitate this, we associate the opacities of the blocks with the labeled rays.
-Specifically, we use a cross-entropy loss $L_{ce}$, between the opacity and mask labels as follows:
+During optimization, the opacity of the block inside the object region tends to approach 1, while the opacity of the block outside the object region tends to approach 0.
+To facilitate this, we associate the opacities of blocks with the labeled rays and apply a cross-entropy loss $L_{ce}$ between the block opacity and the mask labels, defined as follows:
 ${\mathcal{L}}_{\text{opa}}({\mathcal{R}})=\frac{1}{|{\mathcal{R}}|}\sum_{r\in{\mathcal{R}}}L_{ce}\left(\max_{i\in\mathcal{M}}\tau_{i}(x^{r}),l_{r}\right)~{}.$ (11)
 Here, only points $x^{r}$ inside the blocks are sampled, e.g., $\left\{x^{r}\in{\mathcal{X}}_{r}\mid D_{i}(x^{r})\leq 0\right\}$.
-The total loss is the weighted summation of the loss terms described above:
+The total loss is the weighted sum of the loss terms described above:
 ${\mathcal{L}}={\mathcal{L}}_{\text{ren}}+\lambda_{\text{cov}}{\mathcal{L}}_{\text{cov}}+\lambda_{\text{over}}{\mathcal{L}}_{\text{over}}+\lambda_{\text{par}}{\mathcal{L}}_{\text{par}}+\lambda_{\text{opa}}{\mathcal{L}}_{\text{opa}}.$ (12)
-Adaptive Number of Blocks.
-Given that the number of parts varies across scenes, we allow the number of blocks to adjust during optimization adaptively.
-When the opacity of an existing block falls below a threshold $t$, this block is immediately removed.
-Additionally, we explore a block-adding mechanism using point clouds that are randomly generated within the scene area.
-Specifically, we use DBSCAN [34] to cluster initial point clouds (e.g., accessible by-products of COLMAP [33]) that are not covered by any blocks.
-A new block is then introduced at the center of each point cloud cluster.
-The other parameters of these new blocks are initialized randomly.
-At last, block-level reconstruction gets a 3D scene composited by several superquadrics, as shown in Fig. 2.
-3.2.2 Point-level Optimization
-The primitive-based hybrid representation effectively decomposes the shape into parts but performs suboptimally for complex objects.
-To address this, we introduce a refinement stage for more precise geometry fitting.
-In this stage, we decouple the constraint between Gaussians and superquadrics, enabling the Gaussians to be freely optimized.
-Furthermore, to prevent Gaussian components from one block passing through another and disrupting the continuity and plausibility of the part decomposition, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
+3.2.2 Adaptive Number of Blocks.
+Given that the number of parts may vary across different scenes, we allow the number of blocks to adjust during optimization adaptively.
+Specifically, when the opacity of an existing block falls below a threshold $t$, this block is removed immediately.
+Furthermore, we explore a block-adding mechanism to dynamically incorporate new components, ensuring comprehensive coverage of target objects.
+Specifically, DBSCAN [45] is employed to cluster initial point clouds that are not covered by any blocks, where the point clouds are either randomly initialized or derived as by-products from COLMAP [44].
+For each point cloud cluster, a new block is introduced at its center, with the remaining parameters of the block initialized randomly.
+The final block-level reconstruction generates a textured 3D scene represented by multiple superquadrics, with surface details rendered through Gaussians, as illustrated in Fig. 2.
+3.3 Point-level Decomposition
+The primitive-based hybrid representation effectively decomposes the shape into parts but exhibits suboptimally for complex objects.
+To address this limitation, we perform a refinement stage that enhances geometric fitting.
+In this stage, the constraint between Gaussians and superquadrics is decoupled, enabling independent Gaussian optimization.
+Furthermore, to prevent Gaussian components from one block passing through adjacent blocks and disrupting the continuity and plausibility of the part decomposition, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
 ${\mathcal{L}}_{\text{enter}}=\frac{1}{N}\sum_{x\in\Omega}\sum_{m\in\mathcal{M}\setminus\{\delta\}}\text{ReLU}(-D_{m}(x)),$ (13)
@@
 As seen in Fig.
-2, the reconstruction becomes more accurate and aligns better with the target shape after this stage.
+2, the reconstruction becomes more accurate and aligns better with the target shape.
 Figure 3: Qualitative comparison on DTU [23] and ShapeNet [23] .
@@
 Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
-Figure 4: Qualitative results on BlenderMVS [45] and self-capatured data.
+Figure 4: Qualitative results on BlendedMVS [63] and self-captured data.
 We present RGB renderings and decomposed parts from novel views.
-The top examples are from the BlenderMVS dataset, and the last example is from our captured scenes.
+The top examples are from the BlendedMVS dataset, and the last example is from our captured scenes.
 Figure 5: Ablation studies on key strategies.
@@
 MBF [42] 3D GT 3.12 2.66 3.84 2.54 1.59 2.11 2.19 2.01 2.32 2.45 2.17 2.12 3.83 2.02 2.55 2.50 34.1
-EMS [22] + Neus [40] Iamge 5.99 5.56 4.43 4.32 5.42 6.14 3.75 3.96 4.63 4.34 5.88 5.11 4.29 4.83 3.53 4.97 8.87
+EMS [32] + Neus [55] Image 5.99 5.56 4.43 4.32 5.42 6.14 3.75 3.96 4.63 4.34 5.88 5.11 4.29 4.83 3.53 4.97 8.87
 MBF [42] + Neus [55] Image 2.69 3.37 3.22 2.69 3.63 2.60 2.59 3.13 2.85 2.51 2.45 3.72 2.24 2.49 2.52 2.85 46.7
@@
 Ours (Block-level) Image 5.68 4.91 1.85 2.61 3.75 4.66 3.75 7.57 4.27 4.38 3.49 4.48 3.61 4.21 3.70 4.19 5.9
-Ours (Point-level) Image 0.95 1.38 0.55 0.71 1.02 1.39 1.16 1.33 1.49 0.80 0.95 1.31 0.69 1.20 0.82 1.05 5.9
+Ours (Point-level) Image 0.70 1.17 0.55 0.65 1.06 1.23 1.10 1.36 1.37 0.78 0.92 1.41 0.69 1.05 0.71 0.98 5.9
 Table 1: Quantitative comparison on DTU [23].
@@
 Ours (Block-level) 4.19 19.84 0.82 0.189 $\sim$ 30 m
-Ours (Point-level) 0.95 35.04 0.99 0.015 $\sim$ 40 m
+Ours (Point-level) 0.98 35.04 0.99 0.015 $\sim$ 40 m
 Table 2: Quantitative results on DTU [23].
@@
 4 Experiments
-We evaluate our approach on 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time, comparing it with previous state-of-the-art methods.
+We comprehensively evaluate our approach across four key aspects: 3D reconstruction quality, view synthesis performance, shape parsimony, and computational efficiency, comparing it with state-of-the-art methods in the field.
 4.1 Datasets
-We conduct comparative experiments on two public datasets: DTU [17] and ShapeNet [6].
-To validate the effectiveness of our method in man-made objects, we construct a ShapeNet subset with four categories: Chair, Table, Gun, and Airplane.
-The DTU dataset comprises 15 commonly used scenes, and each category in the ShapeNet dataset contains 15 objects.
-Additionally, to demonstrate our method’s potential with real-world data, more qualitative results are presented on BlenderMVS dataset [45] and self-captured scenes.
+Extensive experiments are conducted on two widely-used public datasets: DTU [23] and ShapeNet [6].
+Specifically, 15 standard scenes from the DTU dataset that are widely adopted for reconstruction quality assessment are used for evaluation.
+Each scene contains either 49 or 64 images.
+Additionally, we construct a ShapeNet subset consisting of four categories: Chair, Table, Gun, and Airplane, with 15 objects selected per category.
+For each object, 50 rendered images are generated for training.
+We further validate real-world applicability on both the BlendedMVS dataset [63] and self-captured scenes, demonstrating robust performance across synthetic and natural environments.
 4.2 Implementation Details
 We build upon the 2DGS [21] architecture and add a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq.
-14.
+13.
 In the ShapeNet benchmark, we observe a common degradation in mesh reconstruction as shown in Fig.
@@
 4.4.
-The same hyperparameters are used for all experiments.
 We set the initial number of primitives $M$ to 8.
@@
 During block-level optimization, the block-adding operation is carried out at the 5k-th and 10k-th iterations.
-The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.01, and 0.01, respectively.
-For more details on the dataset, metrics, baselines, and method, please refer to the supplementary materials.
+The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.002, and 0.01, respectively.
+For more details about datasets, metrics, and baselines, please refer to Supplementary.
 4.3 Evaluations
@@
 3, we compare our geometry reconstruction to SOTA shape decomposition methods on Chamfer distance and training time using DTU and Shapenet dataset.
+The Chamfer distance metrics reported for ShapeNet are scaled by a factor of 100 for readability.
 Our method consistently outperforms in all scenes compared to prior works.
@@
 2 and Tab.
-4, our approach achieves competitive results with advanced Gaussian-based methods(e.g., 2DGS [16]) and surface reconstruction approaches (e.g., Neuralangelo [21]), which are limited to producing unstructured meshes.
-Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than PartNeRF and over 3X faster than DBW.
+4, our approach achieves competitive results with the advanced Gaussian-based 2DGS [21] and surface reconstruction Neuralangelo [30], which are limited to producing unstructured meshes.
+Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than PartNeRF [48] and over 3X faster than DBW [36].
 Appearance Reconstruction.
-Beyond part decomposition, our method enables high-fidelity image synthesis, attributed to integrating Gaussian splatting within superquadric representations.
-EMS and MBF lack image rendering capabilities, while PartNeRF and DBW yield low-quality view synthesis.
+In addition to part decomposition, our method enables high-fidelity image synthesis, attributed to integrating Gaussian splatting within hybrid representations.
+EMS [32] and MBF [42] operate directly on point clouds and consequently lack image rendering capabilities, while PartNeRF [48] and DBW [36] yield low-quality view synthesis.
 In contrast, our method achieves high-quality appearance rendering results, as demonstrated in Tab.
@@
 4.
-Figure 6: Qualitative comparison with SAM-based methods. Our method achieves finer-grained and more distinct part decomposition.
+Figure 6: Qualitative comparison with SAM-based methods.
+Visual results demonstrate that our method produces more structurally coherent decompositions, whereas SAM-based approaches frequently exhibit spatial discontinuities.
 Block-level Point-level
@@
 4.3.2 Real-life Data
-To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from BlenderMVS dataset [45] and a self-captured dataset.
+To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from the BlendedMVS dataset [63] and a self-captured dataset.
 As shown in Fig.
-4, our approach can robustly produce both realistic appearances and meaningful 3D decompositions across a variety of data types.
+12, our approach can robustly produce both realistic appearances and reasonable 3D decompositions across a variety of data types.
 More results are provided in the supplementary material.
@@
 Part-Aware Editing (top): After optimization, we can easily edit the scene by adding, scaling, or moving specific parts.
-3D Content Generation (middle): By combining parts of different objects, we can create new 3D content.
-Physics simulations (bottom): The representation enables us to perform physics-based simulations easily.
+3D Content Generation (bottom): By combining parts of different objects, we can create new 3D content.
 4.3.3 Compared to SAM-based Methods
-We also conduct comparisons to SAM-based methods for this task, as shown in Fig.
+We also conduct comparisons to SAM-based [27] methods for this task, as shown in Fig.
 6.
-Despite impressive advances in 3D segmentation and editing [5, 19, 46, 47] by distilling 2D information, achieving finer-grained semantic disentanglement remains challenging due to indistinct textures and severe cross-view 2D inconsistencies.
+Despite impressive advances in 3D segmentation and editing [5, 26, 65, 66] by distilling 2D information, achieving semantic disentanglement but remains challenging due to indistinct textures and severe cross-view 2D inconsistencies.
 Lifting 2D segmentation features to 3D segmentation will also lead to feature misalignment, leading to visual artifacts.
-In contrast, our method segments directly from 3D space, enabling more fine-grained and clearly defined component decomposition.
+In contrast, our method performs direct 3D space segmentation, yielding more reasonable and structurally coherent decompositions.
 4.4 Ablations
@@
 In this section, we first analyze the design choices by isolating each strategy to assess its impact.
-We report the averaged quantitative performance on 15 instances of chair category in Tab.
+We report the averaged quantitative performance on 15 instances of the chair category in Tab.
 5 and provide visual comparisons in Fig.
 5.
-Removing overlapping loss enhances reconstruction quality but reduces semantic clarity.
-Omitting opacity loss drastically affects the number of primitives and doubles the CD, highlighting its role in controlling primitive presence and reconstruction accuracy.
+Removing overlapping loss provides better reconstruction accuracy but more overlapping parts.
+The opacity loss drastically affects the number of primitives and doubles the CD, highlighting its role in controlling primitive presence and reconstruction quality.
 Coverage loss is crucial in reconstruction accuracy, ensuring primitives align correctly with the target.
 The adaptive primitive strategy fills in missing parts and improves the reconstruction metric.
-Without parsimony loss, CD and rendering metrics are optimal but result in redundancy and over-decomposition.
+Without parsimony loss, CD and rendering metrics are optimal but result in over-decomposition.
+Note that the full model did not yield the highest accuracy because the part-aware reconstruction trades off the accuracy and rationality of the part decomposition.
+The objective is to achieve high accuracy while preserving reasonable parts.
 We also validate the effectiveness of Gaussian scale regularization in Fig.
 7.
-2DGS [16] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF.
-Holes occur since large-scale Gaussians can’t support consistent depth rendering from different views.
+2DGS [21] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF integration [71].
+The presence of holes stems from the inability of large-scale Gaussians to maintain view-consistent depth.
 Scale regularization effectively addresses this issue by suppressing large-scale Gaussians.
@@
 Fig.
-9 illustrates three applications of our method that original 3DGS and NeRF-based approaches do not support.
-Firstly, after optimization, we obtain the part decomposition, facilitating easy editing of specific object or scene components, e.g., adding, moving, removing, or scaling.
+9 illustrates two applications of our method that the original 3DGS and NeRF-based approaches do not support.
+Firstly, after optimization, we obtain the part decomposition, facilitating easy editing of specific objects or scene components, e.g., adding, moving, removing, or scaling.
 Secondly, by combining parts of different objects, our method enables the creation of new high-quality 3D content.
-Additionally, our method enables easy manipulation of objects within a scene, supporting physics-based simulations.
-5 Conclusion
-We introduce a hybrid representation of superquadrics and 2D Gaussians, to learn 3D scenes in a part-aware representation.
-Compared to prior works, the proposed method also keeps geometry details instead of shape abstractions, supporting high-level image rendering quality.
-The method gets state-of-the-art performance in comprehensive evaluations.
-One limitation is that the method takes background-free scenes as inputs, from segmentation tools such as SAM, as we care about the structure of the centered scene the most.
-In the future, we aim to explore how to model backgrounds and extend the work to address entire complex scenes.
+5 Conclusions
+We introduce PartGS, a hybrid representation of superquadrics and 2D Gaussians, to learn 3D scenes in a part-aware representation.
+Compared to prior works, the proposed method retains geometry details, supporting high-quality image rendering.
+It obtains state-of-the-art performance in comprehensive evaluations.
+One limitation is that it takes background-free scene images as inputs, leveraging segmentation tools such as SAM [27].
+In the future, we aim to explore how to model backgrounds and extend to larger and more complex scenes.
+Acknowledgements
+This work is supported in part by the NSFC (62325211, 62132021, 62372457), the Major Program of Xiangjiang Laboratory (23XJ01009), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Natural Science Foundation of Hunan Province of China (2022RC1104).
+Supplementary Material
+At first, this supplementary material provides detailed experimental settings, including data processing procedures, implementation details for comparison baseline methods, and evaluation metrics (Sec.6).
+Subsequently, we provide further implementation specifics, including the training configurations and rendering process (Sec.8).
+In Fig.
+10 and Fig.
+11, we provide additional visual comparisons of our method against state-of-the-art baselines on DTU [23] and ShapeNet [6] datasets.
+Moreover, we include visual results on several scenes from the BlendedMVS [63] and self-captured data in Fig.
+12.
+It can be observed that the decomposed reconstructions by our method are more reasonable and precise, which is believed to be beneficial for downstream tasks.
+Our code and data are available at https://github.com/zhirui-gao/PartGS.
+6 More Details on Experiments
+6.1 Datasets
+We conduct evaluations on two public datasets: DTU [23] and ShapeNet [6].
+DTU is a multi-view stereo (MVS) dataset comprising 80 forward-facing scenes, each captured with 49 to 64 images.
+In experiments, we use 15 publicly recognized scenes commonly adopted in previous studies [21, 30].
+We downsample the images to a resolution of $400\times 300$ for computational efficiency.
+To validate the effectiveness of our method in the decomposition of man-made objects, we construct a subset of the ShapeNet dataset comprising four categories: Chair, Table, Gun, and Airplane.
+Each category includes 15 distinct objects, providing diverse instances.
+For each object, we randomly sample 100 camera poses on a sphere and render images at a resolution of $400\times 400$.
+The rendered images are equally split into training and testing sets.
+Additionally, to explore the potential of PartGS in handling real-life data, we present qualitative results on the BlendedMVS dataset [63] and self-captured scenes.
+For BlendedMVS, we use official camera poses, while for self-captured scenes, camera poses are estimated by COLMAP [44] and normalized using IDR [64].
+All images are resized to $400\times 300$ pixels.
+For the DTU, ShapeNet, and BlendedMVS datasets, we utilize the ground-truth foreground masks provided within the datasets.
+For self-captured real-world scenes, we employ the Segment Anything Model (SAM) [27] to segment foreground objects.
+6.2 Implementation Details on Baselines
+We compare our method with four state-of-the-art works on 3D shape decomposition: EMS [32], MonteboxFinder (MBF) [42], PartNeRF [48] and DBW [36].
+MBF and EMS are applied to point clouds, utilizing cuboids and superquadrics, respectively, to fit the 3D points.
+The input point clouds are sampled from either ground truth 3D shapes (GT) or meshes reconstructed by state-of-the-art MVS method Neus [55].
+We adhere to the testing procedure outlined by DBW [36] and randomly sample 5K and 200K points from the GT meshes as the input for EMS and MBF, respectively.
+Similar to our problem setting, PartNeRF and DBW use images as input to construct structural 3D representations.
+To ensure fair comparisons, we discard the pre-defined ground plane in DBW, as the test scenario consists exclusively of foreground objects.
+For PartNeRF, we perform instance-specific training and set the same number of parts as in our method, while retaining all other default parameter settings.
+Additionally, we compare with SOTA Gaussian Splatting reconstruction method 2DGS [21], and surface reconstruction approach Neuralangelo [30], to provide an intuitive evaluation of the reconstruction quality achieved by our method.
+6.3 Evaluation Metrics
+We evaluate from 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time.
+  • The 3D reconstruction quality is measured by the official Chamfer distance (CD) evaluation [23] between the recovered geometry and GT, reflecting the accuracy of 3D reconstructions.
+  • View synthesis uses three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [57], and Learned Perceptual Image Patch Similarity (LPIPS) [69].
+  • Shape parsimony is quantified by the number of parts, while reconstruction time refers to the running time measured on the same device.
+While shape parsimony does not directly indicate the quality of the decomposition, a smaller number of decomposed components, given consistent reconstruction quality, suggests a more concise and reasonable decomposition.
+This is because fitting different parts of an object with fewer blocks is inherently more challenging, whereas utilizing more blocks simplifies the fitting process.
+Figure 10: Qualitative comparisons on DTU [23].
+We compare our approach with state-of-the-art baselines on the DTU dataset with the background removed.
+The last two columns show our block-level and point-level reconstructions, respectively.
+Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
+Figure 11: Qualitative comparisons on ShapeNet [23].
+We compared our approach with state-of-the-art baselines across four categories.
+The last two columns display our block-level and point-level reconstructions, respectively.
+Our method uniquely provides reasonable 3D part decomposition while simultaneously capturing detailed geometry.
+Figure 12: Qualitative results on BlendedMVS [63] and self-capatured data.
+We demonstrate the RGB renderings and decomposed parts from novel views across a variety of objects.
+The first five examples are from the BlendedMVS dataset, and the remaining examples are from our own captured scenes.
+7 Additional Comparative Experiments
+Very recently, DPA-Net [68] and GaussianBlock [24] have achieved advanced 3D shape abstraction.
+DPA-Net enables part-aware reconstruction in a feedforward manner, requiring approximately 3 days of GPU training on a pre-collected dataset and 2 hours per object for inference.
+GaussianBlock utilizes SAM [27] to guide superquadric splitting and fusion for 3D decomposition, with a processing time of 6 hours per object.
+We compare our method on DTU and BlendedMVS scenes, with results presented in Fig.
+13.
+The proposed approach achieves comparable performance while demonstrating superior efficiency.
+It is important to clarify the supervision requirements: DPA-Net is a supervised method that relies on SAM-generated segmentation masks for training, whereas GaussianBlock, while not strictly 3D supervised, depends on high-quality pre-trained datasets that typically require days of training.
+Our method distinguishes itself by being completely self-supervised, requiring neither 3D supervision nor pre-trained components, yet achieving significantly better computational efficiency.
+Figure 13: Qualitative comparisons to DPA-Net and GaussianBlock. The first two examples are from the DTU dataset, and the last two examples are from the BlendedMVS dataset.
+8 Implementation Details
+8.1 Training Configurations
+To ensure uniform distributions of 2D Gaussians across surfaces of the mesh, random barycentric coordinates are generated directly within each triangular face.
+Specifically, barycentric weights are computed as $u=\sqrt{rand}$ and $v=rand$.
+These weights are transformed to obtain $\alpha=[1-u,u(1-v),uv]$, $rand\in(0,1)$.
+The sampled position on triangular face is calculated as $o=\alpha_{0}v0+\alpha_{1}v1+\alpha_{2}v2$, where $v_{o},v_{1},v_{2}$ are vertices of the triangle.
+This method compensates for the non-uniformity caused by the triangle’s geometry.
+Unlike naive random sampling which results in uneven distributions, it ensures that the sampled points are evenly distributed across mesh surfaces.
+The same hyperparameters are used for all experiments.
+We set the initial number of primitives $M$ to 8.
+In the hybrid representation, each superquadric mesh is a level-2 icosphere (320 triangular face).
+Each triangular face contains 100 Gaussians, with a scaling parameter $c$ of 0.1.
+The number of sampled points in each ray is 2048.
+During refinement, we employ regularization in 2DGS to achieve better geometric reconstruction, including depth distortion maps, depth maps, and normal maps.
+Additionally, we introduce a mask cross-entropy to filter out extra noise Gaussians.
+To extract the meshes from 2D Gaussians, we use truncated signed distance fusion (TSDF) to fuse rendered depth maps, utilizing Open3D [71].
+8.2 Rendering
+With Gaussians attached to the surface of each block, we achieve view-dependent rendering through tile-based rasterization as Gaussian Splatting [25, 21].
+Given a view, Gaussians on the superquadric surface are projected onto the image space, forming an RGB image.
+Initially, the screen space determines the bounding box for each Gaussian.
+Subsequently, these Gaussian ellipses are sorted according to their depths of center to the image plane.
+Finally, volumetric alpha compositing [41] is utilized to integrate the alpha-weighted RGB values for each pixel.
+To formulate the process, considering a viewing transformation ${W}$, the covariance matrix ${\Sigma}_{i}^{{}^{\prime}}$ of $i$-th Gaussian in the camera coordinate system is calculated by:
+${\Sigma}_{i}^{{}^{\prime}}={J}{W}{\Sigma}{W}^{T}{J}^{T},$ (14)
+where ${J}$ is the Jacobian of the affine approximation of the projective transformation, and ${\Sigma}$ is the covariance matrix of the Gaussian ellipse.
+Note that the last row and column of ${\Sigma}$ are omitted since we adopt 2D Gaussians.
+Following alpha compositing, we first calculate an alpha value for each Gaussian ellipse:
+$\alpha_{i}=\tau_{i}\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{u}_{i})^{T}{\Sigma_{i}^{{}^{\prime}}}_{-1}(\mathbf{x}-\mathbf{u}_{i}))).$ (15)
+Here, $\mathbf{u}_{i}$ is the center coordinate of the projected Gaussian ellipse, and $\tau_{i}$ is the opacity of the block where the $i$-th Gaussian is located.
+The calculated alphas are sorted according to their depths from the image plane.
+Meanwhile, we can acquire the color value $c_{i}$ from the spherical harmonics of $\mathcal{N}$ ordered points, thus obtaining the rendering RGB value:
+$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (16)
+This process is differentiable and can optimize the hybrid representation through gradient descent.